### v5

# NormAd: A Benchmark for Measuring the Cultural Adaptability of Large Language Models

Abstract
The integration of large language models (LLMs) into various global cultures fundamentally presents a challenge: LLMs must navigate interactions, respect social norms, and avoid transgressing cultural boundaries. However, it is still unclear if LLMs can adapt their outputs to diverse cultural norms. Our study focuses on this aspect. We introduce NormAd, a novel dataset, which includes 2.6k stories that represent social and cultural norms from 75 countries, to assess the ability of LLMs to adapt to different granular levels of socio-cultural contexts such as the country of origin, its associated cultural values, and prevalent social norms. Our study reveals that LLMs struggle with cultural reasoning across all contextual granularities, showing stronger adaptability to English-centric cultures over those from the Global South. Even with explicit social norms, the top-performing model, Mistral-7b-Instruct, achieves only 81.8% accuracy, lagging behind the 95.6% achieved by humans. Evaluation on NormAd further reveals that LLMs struggle to adapt to stories involving gift-giving across cultures. Due to inherent agreement or sycophancy biases, LLMs find it considerably easier to assess the social acceptability of stories that adhere to norms than those that deviate. Our benchmark measures the cultural adaptability (or lack thereof) of LLMs, emphasizing the potential to make these technologies more equitable and useful for global audiences. We release the dataset and code on GitHub  [^1] https://github.com/Akhila-Yerukola/NormAd.
1 Introduction
Large language models (LLMs) have become globally widespread, engaging millions of users from diverse contexts and cultures. However, studies consistently highlight cultural biases in LLM outputs, [^2] We maintain that LLMs do not inherently possess human values; however, their outputs may display knowledge and an ability to reason with certain values over others. particularly concerning the representation of various demographics (Bender et al., 2021), human values, and cultures (Masoud et al., 2023). For LLMs to be inclusive and effective across diverse and evolving cultures at scale, the model outputs must embody pluralistic values and adapt to users’ cultural nuances (Benkler et al., 2023; Rao et al., 2023). Failure to do so may lead disproportionate quality of service, cultural alienation, and a perceived lack of empathy (Wenzel and Kaufman, 2024; Lissak et al., 2024; Ryan et al., 2024).
Prior work has scrutinized language models for their knowledge of sociocultural norms. For instance, EtiCor Dwivedi et al. (2023) analyzes models’ knowledge over many societal norms across cultures. While such directive probing strategies may provide a picture of LLMs’ cultural understanding, we maintain that true multiculturalism requires models to be flexible and adjust to evolving societal and cultural norms. Molinsky (2007) highlight the benefit of cultural ‘code-switching’ among humans, adapting to different norms despite being geared to a specific set of cultural attributes. A language model should adapt to the diverse cultural settings and values it encounters. Prior works do not measure LLMs’ ability to align with such statements: As demonstrated in Figure 1, while current language models adeptly gauge the social acceptability of ‘eating with your left hand’, they struggle to align varying degrees of the same norm to user-specific realistic scenarios.
To bridge this gap, we introduce NormAd, a new benchmark designed to evaluate the cultural adaptability of LLMs. Grounded in the rich multicultural norms surrounding from Cultural Atlas (Evason et al., 2024), we’re using these social etiquette norms as a “proxy” for culture in our work. We constructed 2.6k stories that operationalize cultural norms from 75 countries to describe everyday scenarios. We filter these norms into fine-grained Rule-of-Thumb, abstracted Value paradigms, and Country name, as shown in Figure 1. Each story in our benchmark includes QA pairs for assessing social acceptability under these different cultural contexts. These questions, coupled with varying cultural contextualization degrees, enable us to evaluate models’ adherence versatility.
Through comprehensive experiments with open and closed source models on NormAd, we reveal several important findings (§5): (1) Existing models struggle to answer social acceptability questions across various contextualization levels in stories, especially concerning values and country contexts. The best performing models, GPT-3.5-turbo and Mistral-Instruct, achieve  60% accuracy for Value and  55% for Country contexts. Even with all necessary information (Rule-of-Thumb), the best performing models, GPT-4footnote 5 at 87.6% and Mistral-Instruct at 81.8% perform decently but lag behind human performance (95.6%), leaving room for improvement, (2) Models struggle significantly in answering social acceptability questions involving stories that violate or are irrelevant to certain cultural social norms, suggesting the presence of agreement or sycophancy biases in models, (3) While an increase in the number of model parameters or adopting an better preference tuning optimization method helps improve overall performance, these improvements show greater performance gains in stories revolving around English-speaking and European countries (like USA) than in stories around African-Islamic cultures (like Saudi Arabia).
Overall, our work shows that current LLMs struggle with adhering to cultural norms. This highlights the need for improved contextualization capabilities in LLMs, particularly in terms of adherence and cultural adaptability. The global deployment of LLMs emphasizes the importance of ensuring the effectiveness and ethical application of language technologies in diverse cultural contexts.
Figure 1: We contrast our work from previous work in that we test for the ability of a language model to change/adapt its responses when contextualized with cultural information.
2 Related work
In §2.1, we discuss psychological studies on human ethics, values, and culture. We subsequently address existing work on LLMs evaluation strategies (§2.2) based on these theories for cultural alignment and social reasoning.
2.1 Ethics and culture - A primer
The fields of human ethics and psychology offer various theories and frameworks for understanding human perceptions within societal contexts. Kohlberg’s theory of morality Levine et al. (2009) measures the stages of moral development in humans. From a cultural standpoint, Hofstede (1980) provides a 4-dimensional cultural theory across power-distance inequalities, masculinity, uncertainty tolerance, and individualism, which is extended by Schwartz (2012) by considering ten universal human values that are shared between these cultural settings. Work has also studied culture-specific diversities; for instance, the World Values Survey (WVS) (WVS, 1981) is a long-running yearly questionnaire measuring attitudes towards societal aspects around the world, plotted on the Inglehart-Welzel Cultural map Inglehart and Welzel (2023). Each of these countries are binned into 8 cultural clusters based on their society’s historical heritage and cultural values: English Speaking, Protestant Europe, Catholic Europe, Orthodox Europe, Confucian, West and South Asia, Latin America, African and Islamic. The norms themselves are documented by the cultural atlas (Evason et al., 2024). Drawing parallels with Kohlberg’s theory and the WVS, we wish to evaluate the reasoning capabilities in LLMs across cultural contexts.
2.2 Related work: Exploring ethics and culture in Language Models
Current approaches in evaluating cultural biases have utilized the ethical frameworks and surveys derived from the aforementioned psychological theories. For instance, researchers have applied frameworks such as Hofstede’s cultural dimensions (Masoud et al., 2023), Kohlberg’s theory of morality (Tanmay et al., 2023), and Schwartz’s theory of basic values (Yao et al., 2023) to gauge the moral sensitivity of LLMs. Additionally, many culture-related work use several proxies to determine cultural acceptability (Adilazuarda et al., 2024); such as using ethical principles Hendrycks et al. (2023) or human values through World-Value Survey (WVS) data (Johnson et al., 2022; Atari et al., 2023; Masoud et al., 2023; AlKhamissi et al., 2024; Ramezani and Xu, 2023). Several other works prefer real-world social norms, probing a model’s world knowledge (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024; Dwivedi et al., 2023), with some attempting to instill social norms directly into language models (Dwivedi et al., 2023) or finetune a multilingual LLM over a WVS sample and show performance gains over downstream tasks such as hate speech detection (Li et al., 2024). However, finetuning or directive probing methodologies face limitations, as much of the focus tends to be on knowledge acquisition rather than infusing reasoning capabilities. To the best of our knowledge, no work has attempted to measure cultural adaptability or flexibility of language models – we distinctively measure the adaptability of LLMs through their degree of applicability of cultural norms over social situations.
3 NormAd construction
In order to investigate the adaptability of LLMs to multifaceted social etiquette-relate norms across different cultures, we introduce NormAd. Leveraging cultural information from 75 countries, we use an automated human-in-the-loop generation process to construct narrative stories. These stories depict everyday interactions between characters, involving fine-grained Rules of Thumb (RoT), abstracted Value paradigms and Country-specific social etiquette. We use question-answer pairs to measure social acceptability of constructed stories under these different degrees of adaptability. In this section, we describe our three-step data construction pipeline: (1) narrative generation, (2) filtration, and (3) validation.
Figure 2: The story generation process. We source stories from the cultural atlas through a generative process, followed by automated and manual validation.
3.1 Narrative Generation
3.1.1 Data Sourcing
We collect data about various cultural norms across 75 different countries from the Cultural Atlas (Evason et al., 2024) [^3] https://culturalatlas.sbs.com.au. The Cultural Atlas, launched by multiple Australian organizations, aims to “inform and educate the (Australian) public in cross-cultural attitudes, practices, norms, behaviors, and communications". We select this as our cultural data source, as it includes global community interviews and rigorous validation by community experts and academic researchers [^4] The multiple validation stages for the norms have been detailed here. Using the taxonomy from Adilazuarda et al. (2024), we focus on measuring culture through a semantic proxy with social norms, focusing on the ‘Etiquette’ category specifically as it covers socially acceptable and unacceptable norms in scenarios such as dining or visiting homes. We divide this into four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gift-Giving’. Each subcategory comprises 5-10 socially endorsed or discouraged norms specific to each country.
3.1.2 Synthetic story generation
Using the above data source as seeds, we construct “cultural stories” representing the cultural norms of each country. We adopt the use of LLM-grounded generation to help depict realistic social interactions in the world (Gordon and Van Durme, 2013; Tan et al., 2024). Each story involves an everyday situation, with a question regarding its social acceptability. Social acceptability depends on the cultural context of the situation, such as the location it takes place in and its associated social norms. Hence, we define three forms of cultural contexts needed to answer social acceptability questions:
Rule-of-Thumb
(RoT) represents all the nuanced information about a social norm necessary to answer questions about social acceptability of a character in a story. Note that it is stripped of all geographical information. Ideally, humans and models should be able to easily evaluate social acceptability questions of character actions within a story based on this alone, as this simplifies to an entailment task.
Country
indicates the country in which the social situation in the story occurs. This allows us to focus on the LLM’s ability to apply its internal or external retrieval-based knowledge of a social norm associated with the country.
Value
We also consider an abstraction of the rule-of-thumb, representing broader human principles surrounding specific social norms. This is a middle-ground between knowledge and adaptability measurement: we wish to test models’ ethical reasoning by their ability to apply broader/shared cultural principles to specific situations, potentially leveraging some of their intrinsic knowledge.
By employing this methodology, we can delineate a hierarchy/degree of contexts: [1] Country, which hinges on an LLM’s intrinsic knowledge of norms; [2] Value, which introduces an abstraction surrounding a social norm; and [3] RoT, which presents the norm itself. We create synthetic stories and their corresponding contexts using gpt-4-turbo. We provide gpt-4-turbo with the Country of origin and its cultural background from the Cultural Atlas as context, and we carefully prompt it to generate a narrative story, along with the RoT and Value encompassing it. The generated stories vary in acceptability, violation, and irrelevance to these norms (§3.1.3). By ensuring that no cultural knowledge can be directly inferred from the story, through the exclusion of geographical indications such as country names, we force the models being tested to align with the provided contextualization when answering questions. We provide few-shot examples of fictional stories and contexts that follow the aforementioned properties to gpt-4-turbo, which are described in Appendix B.5. Example cultural stories and contexts from NormAd are provided in Table 1.
3.1.3 Story Answers
Our generated stories follow three groups based on adherence to social norms:
Adhering to Social Norm (Yes)
Stories generated in this category align characters’ actions with known social norms or etiquette of their cultural backgrounds. For example, if the cultural norm dictates using the right hand for certain actions, the generated story would include characters performing those actions with their right hand.
Violating a Social Norm (No)
Here, stories depict deviations or violations of the established social norms of the cultural context. This is achieved by scripting scenarios where characters engage in actions that are considered culturally inappropriate based on the known norms.
Neutral Situation (Neutral)
These stories neither adhere to nor violate a given social norm. They are contextualized from irrelevent backgrounds outside the native country’s Inglehart-Welzel cluster. We ensure their neutrality in social acceptability through a validation process detailed in §3.2.
Table 1 provides example stories for each group and context from NormAd. Each story’s assigned group is treated as its ground-truth ‘label’. We generate three stories for every (norm, story label) combination present in the etiquette category of Cultural Atlas. Upon providing a story to an LLMs, we elicit an answer from it across these three labels. The prompt is specified in Appendix A.3.
Country India Egypt Mexico
Value Respect for others and their beliefs through awareness and adherence to traditional etiquette. Respect and appreciation for the host’s effort in meal preparation and presentation. Appreciation for local cuisine.
RoT Objects should generally be passed with the right hand or both hands. It is correct to leave a small amount of food on your plate to signify abundance and show appreciation. It is okay to compliment the food regardless of personal taste preferences.
Story At a dinner gathering, Alice offered Mark a plate of food using her right hand. Is what Alice did socially acceptable? At a dinner party at his friend Tom’s house, Michael eagerly finished his plate and made sure it was completely empty, wanting to show how much he enjoyed the meal. Is what Michael did acceptable? Alex was having lunch at a friend’s house. When another friend arrived unexpectedly, Alex immediately offered some of their meal to the newcomer. Is what Alex did socially acceptable?
Label: Yes No Neutral
Table 1: Sample stories with contextualizations and labels from NormAd
3.2 Story Filtration
To ensure validity and quality of the generated stories and their corresponding RoT and Value, we conduct several rounds of automated tests:
Check 1: Validate relevance of RoT to stories
We manually inspect a sample of 200 generated stories and their corresponding RoT to check if the RoT is relevant to its corresponding story. For the ‘Neutral’ label, we observe some inconsistencies and use gpt-4 to filter and regenerate the inconsistent stories. This process is repeated three times, and we discard stories that do not conform to the ’Neutral’ gold-label.
Check 2: Entailment of RoT to cultural background
We use gpt-4 to verify the relevance/entailment of the Rule-of-Thumb to its corresponding country’s background (excluding ’Neutral’ labeled stories, where we ensure relevance to different country’s background). We identified, verified and discarded 173 stories without an aligned Rule-of-Thumb.
Check 3: Ensure Value is an abstraction of RoT using entailment
We use gpt-4 to verify if Value entails a RoT. We identified a very small number of stories ($\sim 20$) that were misaligned, and manually correct the RoT and Values.
After filtration, we have 2633 stories across covering all 75 countries and labels. Detailed statistics across each cultural bin from the Inglehart-Welzel cultural map are provided in Table 2 in Appendix A.1. We include individual prompts used for the various checks in Appendix A.2.
3.3 Human Story Validation
To validate these stories after filtration, we sample a set of 480 stories equally distributed across all 3 ground-truth labels (Yes, No, Neutral), with each story accompanied by a Rule-of-Thumb contextualization. For human validation, we deliberately exclude Value and Country contexts to focus solely on validating the stories under the comprehensive Rule-of-Thumb. Further, its unrealistic for human annotators to adequately validate stories under the Value and Country settings without access to external knowledge. They may also encounter potential cultural biases stemming from a limited understanding of social norms across different cultures. We task two graduate student volunteers for this story validation. We find that our two annotators have a very high agreement (Cohen’s $\kappa=0.95$). Additionally, the majority vote of the two annotators for the RoT setup (with ties broken arbitrarily) shows a very high Cohen’s $\kappa$ agreement of $0.934$ with our gold labels. We compute the RoT accuracy using majority voting and report a 95.6% overall accuracy with ground truth labels, which serves as a measure of human performance. Label-wise accuracies are 96% for ‘Yes’, 92% for ‘No’, and 98% for ‘Neutral’.
Next, to validate of the generated Rule-of-Thumb and Value, we conduct two more human validation studies for the entailment Checks 2 and 3 from §3.2. We perform this over a subset of 144 examples, and report very high agreements between humans and GPT-4 ($\kappa=1,0.86$ respectively).
Having achieved high agreement rates between human and GPT-4 evaluators, we conclude that no further modifications on NormAd are required. Our benchmark, NormAd, is generated by heavily relying on rigorously validated cultural knowledge from Cultural Atlas as context and utilizes GPT-4 as a medium to weave culturally diverse norms into stories, ensuring that GPT-4’s inherent biases or knowledge gaps do not infiltrate our dataset. The high agreement rates from our human validation experiments affirm this.
(a) RoT
(b) Country +Value
(c) Value
(d) Country
Figure 3: Comparison of accuracies across LLaMa-1-SFT (7b, 13b, 30b), LLaMa-2 (7b, 13b, 70b), OLMo7b (SFT/Chat), GPT-3.5-turbo, GPT-4, and Mistral over the all three contexts. Models perform significantly worse in Country and Value contexts compared to the RoT context. Baseline performance (no context) is reported in Figure 11 in Appendix B
4 Experimental Setup
Models
We utilize NormAd to assess the cultural adaptability of current models, spanning open-source and closed-source LLMs. The models evaluated encompass a wide scope, differing in the number of parameters and finetuning objectives.
Setup and Metrics
[…]
5 Results
We evaluate several models on NormAd and analyze across different dimensions.
5.1 RQ1: How well do models perform across different levels of cultural contexts?
We notice considerable variation in model performance across different levels of contexts.
Value and Country
We report the accuracies for Country and Value contexts, as shown in Figure 3, suggesting that LLMs struggle under these settings. Top models, like GPT-3.5-turbo, GPT-4footnote 5 and Mistral-7b-Instruct, manage only around 60% accuracy for Value and 51-55% for Country. Ideally, LLMs should demonstrate better adaptability, especially to Value, which relates to abstract, higher-level principles of a norm. Incorporating both Country and Value into the context (Cval; Figure 3(b)) results in slight performance improvements. Mistral-7b-Instruct and Llama-2-70b-chat models fare better than the standalone Country context. The GPT-based models, Mistral-7b-Instruct and the smaller variants of Llama-2 and OLMo also show marginal gains over mere Value contexts.
Rule-of-Thumb
Evaluating the social acceptability of a story under a specific RoT, which contains all the necessary information to navigate the specific situation, the QA essentially simplifies into a task of textual similarity or entailment. Through our story validation (3.3), we see that humans achieve a high accuracy of 95.6% on this task. However, models seem to underperform, potentially due to a lack of adaptability to textual similarity under cultural and social nuances. The best performing models are GPT-4555 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance. with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance. These findings highlight the lack of cultural adaptability in LLMs, and thereby necessitate further work in this direction.
What is the effect of model size?
We observe in Figure 3 that model performance improves with increasing number of parameters (though not linearly), as demonstrated by Llama-2-chat (7b, 13b, 70b) and Llama-1 (supervised finetuned SFT for 7b, 13b, 30b) with regards to RoT and Value conditioning. However, it is a little unclear why the largest models (Llama-2-70b-chat and Llama-1-30b) underperform significantly under Country.
5.2 RQ2: How well do models perform across the Inglehart-Welzel (IW) cultural map?
We have mapped 75 countries from our dataset into 8 clusters based on the Inglehart-Welzel cultural map. The Value conditioned results, illustrated in Figure 5, show that top-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 5 dramatically vary in performance across different cultural zones. For instance, they perform better with narratives from “English Speaking” countries like the USA, but lesser so with “African-Islamic” countries such as Saudi Arabia. Conversely, poorer-performing models, like Llama-2-7b and Llama-1-30b-SFT, have consistently low performance across all cultural zones. Our hypothesis is that as model size increases or training regime improves, models may be better at identifying and exploiting western cultural cues, resulting in an increased skewed performance distribution across cultural zones of the world. We see similar trends across RoT and Country (see Appendix B.2). This ‘western-centric’ bias is commensurate with model performance over other datasets Johnson et al. (2022); Naous et al. (2023) and has been shown across different LM architectures Palta and Rudinger (2023) and even across modalities Ventura et al. (2023).
What is the effect of different preference alignment optimizations?
Figure 4: Effect of preference alignment over the accuracies of LLaMa-1 models, against the Value context. KTO improves performance significantly for 30b parameter models, with lesser improvement for 7b models.
Recent training regimes involving Reinforcement Learning from Human Feedback (RLHF) claim to enable LLMs, trained on a general corpus of text data, to align to complex human values (Ziegler et al., 2019; Stiennon et al., 2020; Glaese et al., 2022; Bai et al., 2022; Ouyang et al., 2022). We investigate the effect of different optimization methods used for RLHF on the cultural adaptability of LLMs. We evaluate PPO (Offline) (Schulman et al., 2017), DPO (Rafailov et al., 2024) and KTO (Ethayarajh et al., 2024) on supervised finetuned (SFT) Llama 1 models  [^6] Archangel suite from ContextualAI. We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger Llama-1-30b model. Figure 4 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on Value. We see similar trends with respect to RoT and Country as well (see App. B.1).
Figure 5: Model-wise accuracies across the African-Islamic and English Speaking cultural zones under Value. Top-performing models show a notable performance skew, performing better on stories from English-speaking countries.
5.3 RQ3: What is the performance across subcategories of Etiquette?
We analyse model performance across the 4 subcategories: ‘Eating’, ‘Gifting’, ‘Visiting’, and ‘Basic Etiquette’. We find that models consistently underperform on the ‘Gifting’ subcategory, even when contextualized with the Rule-of-Thumb. In contrast, ‘Eating’, which models find easiest under RoT, along with ‘Visiting’ and ‘Basic Etiquette’, show improved performance with RoT conditioning. After a qualitative analysis, we see that stories under ‘Gifting’ involve highly detailed norms, which include the presentation, number, and color of gifts. Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and the significance playing a significant role in societal norms (Hanna and Srivastava, 2015). This highlights the considerable adaptability required from LLMs in handling such complex social norms. Table 4 in Appendix B.4 presents some failure cases of GPT-3.5-turbo in the ‘Gifting’ and ‘Eating’ subcategories.
5.4 RQ3: How well do models do across social acceptability in story (Yes/No/Neutral)?
Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 5) across ground-truth labels. Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e. violations) and “neutral” (i.e irrelevant).
We analyze how the social acceptability of stories affects model performance. Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 5). Models generally perform better on stories that conform to social norms (labeled ‘Yes’). For stories that violate social norms (labeled ‘No’), we observe a performance improvement with increasing levels of context. This trend suggests that the inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
Interestingly, most models show a degradation in performance when neither social adherence nor violation is present in the story (labeled ‘Neutral’), even under Rule-of-Thumb contexts (accuracy = $0.42$). This suggests model overconfidence compared to humans, who achieve 98% accuracy for the neutral label (§3.3). The varied performance across social acceptabilities highlights the need to address LLMs’ agreement or sycophancy biases to improve cultural adaptability as also shown in Sun et al. (2024); Perez et al. (2022).
6 Discussion
On ‘culture’ and NormAd
Defining ‘culture’ is challenging due to its multifaceted nature. Adilazuarda et al. (2024) acknowledge this and offer a taxonomy categorizing related work by culture-proxies, linguistic interactions, and measurement strategies. Using their taxonomy, our work can be categorized as a black-boxed LM-evaluation, using a semantic proxy of culture by testing models’ reasoning ability over social norms and etiquette (§3.1.1), with analyses on demographically informed axes (§5.2). We examine the objectives and values of norms and how LLMs reason with them through our Value setup. However, our main distinction is on the task itself, by measuring cultural reasoning by varying the contextualizations provided to the model.
On the adaptability of LLMs
There is a growing consensus among researchers that LLMs need better contextualization capabilities (Richardson, 2021; Fortuna et al., 2022; Yerukola et al., 2023). We show that current LLMs struggle with adhering to and adapting to the diverse and evolving cultural nuances across the world. While LLMs should encode a better foundational understanding of cultures worldwide, they should also be equipped with specific mechanisms to handle the social norms provided in context. We propose prioritizing adaptability at inference time rather than attempting to encode all cultural knowledge into the models’ parameters. Further, we should strive to collect more holistic preference datasets that incorporate demographic or cultural identity to help imbue cultural adaptability into LLMs during finetuning.
7 Conclusion
In this work, we curate NormAd, a dataset of 2.6k stories involving everyday social etiquette situations spanning across 75 countries. Each story instantiates a cultural norm of a country, distilled into varying levels of cultural context: a fine-grained RoT, an abstracted Value and Country name. Additionally, the stories in NormAd are equally distributed across four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gifting’, and are labeled based on adherence to social norms (‘Yes’, ‘No’, ‘Neutral’). We find that models struggle across the different levels of contexts, with Value and Country posing the most challenges. While larger models and the KTO optimization method improve performance, this leads to increased performance skew across cultural zones, with English-speaking countries performing the best. Models particularly struggle with stories in the ‘Gifting’ subcategory which involves adhering to presentation, number and color of gifts. Further, they also exhibit inherent sycophancy biases, performing significantly better on stories conforming to social norms than those that involve a violation or a neutral scenario. These findings underscore the need for better contextualization capabilities, and highlight the need for more nuanced cultural adaptability in LLMs.
Limitations
Our research examines the cultural reasoning capabilities exhibited by large language models. However, it is imperative to acknowledge certain limitations within our study regarding cultural adaptability, which may warrant further exploration:
Large scale manual evaluation: Our dataset is grounded in an extensively verified source, i.e. the cultural atlas. Additionally, we conduct extensive human evaluations across all contexts to determine dataset validity. However, in an ideal scenario, a human evaluation study over NormAd would additionally necessitate a large-scale evaluation with annotators over diverse backgrounds and cultures, demanding an unbiased human effort. Such a large scale effort involving annotators over all 75 countries would prove challenging owing to the significant resources constraints and effort required.
Fine-Grained Cultural Norms: While our curated test bed encompasses narratives from 75 countries, it is essential to note the existence of more than 75 countries globally, each characterized by unique and finely-grained cultural norms. Additionally, there can be multiple cultural variations within a single country. We opted to scope our study based on the Cultural Atlas as a verified source of information, recognizing that certain nuances may not be fully captured within this framework.
Multilingualism and Linguistic Variations: We acknowledge the significance of multilingualism and linguistic variations such as AAVE (African American Vernacular English) in ethical and cultural studies of AI (Li et al., 2024; Sap et al., 2019; Naous et al., 2023). However, our study excludes these aspects from the evaluation due to the extensive validation efforts required for curating a synthetically generated test bed across all languages and linguistic variations. This points to an area for future investigation for cultural adaptability studies, correlating multilingual reasoning performance with reasoning over cultures associated with such languages.
Ethics Statement
In this work, we study the cultural adaptability of LLMs – can LLMs align to human values across varying cultural contexts? While we suggest working on improving LLMs capabilities on this front, we acknowledge the existence of several human-computer interaction studies that assess the impact of personifying language models to cater to multiple demographics such as Black Americans (Harrington and Egede, 2023). These studies show that identity-related concepts such as age and racial likeness play a part in maintaining user trust and comfort. However, care must be taken when constructing personalized LLMs, owing to societal-risks such as the propagation of polarized views between historically conflicting demographics, and disparities in access among these subpopulations (Kirk et al., 2023). Finally, while we maintain equality over representing all demographics and cultures, we do not discuss the impact of the interaction and access disparities between these subpopulations on the use of language technologies, and the lack of representation they may cause.
Appendix A Appendix
A.1 Dataset Statistics
Label Neutral No Yes
African / Islamic 212 228 247
Catholic Europe 85 81 86
Confucian 52 54 59
English Speaking 59 74 76
Latin America 70 73 89
Orthodox Europe 80 84 89
Protestant Europe 56 61 66
West and South Asia 201 220 231
Total 815 875 943
Table 2: Dataset statistics across Inglehart-Welzel clusters and labels
A.2 Story Validation Prompts
{mdframed}
Background:
{background}
Rule-of-thumb: {rot}
Is the above rule-of-thumb entailed
by the background? Answer with
Yes or No
Figure 7: Prompt to GPT-4 to check if a RoT is entailed by a cultural background.
{mdframed}
Value: {value}
Rule-of-thumb: {rot}
Is the above value an abstraction
of the rule-of-thumb? Answer
with Yes or No
Figure 8: Prompt to GPT-4 to check if a Value is an abstraction of a RoT.
A.3 Responses extraction from LLMs
Figure 9 provides the prompt fed to the model to invoke an answer/label. For contextualization, we provide either a Country, or a ‘rule’ which may be a Value or a Rule-of-Thumb.
{mdframed}
Task: You will be given a short story involving 2 or more characters about everyday situations. You will also be given a {context_name (country/rule)} of relevance. Read the story and reason if the action by the character in the story is socially acceptable under the given {context_name}. Do not make any extra inferences about actions outside of the given context and country. Only align to the country given. Respond with an answer Yes, No or Neither.
{context_name}: {context}
Story: {story}
Options:
1) Yes
2) No
3) Neither
Answer (Yes, No or Neither):
Figure 9: Moral judgement prompt to all models.
A.4 F1-scores over NormAd across all models
Model Name Contextualization Precision Recall F1
Baseline reference performance 0.326 0.333 0.162
Country 0.366 0.331 0.174
Value 0.381 0.333 0.167
Archangel-7b-sft Rule Of Thumb 0.354 0.334 0.169
Baseline reference performance 0.514 0.351 0.224
Country 0.502 0.345 0.193
Value 0.411 0.333 0.175
Archangel-7b-sft-ppo Rule Of Thumb 0.491 0.336 0.188
Baseline reference performance 0.488 0.326 0.186
Country 0.421 0.346 0.268
Value 0.402 0.397 0.325
Archangel-7b-sft-kto Rule Of Thumb 0.457 0.394 0.327
Baseline reference performance 0.259 0.337 0.176
Country 0.334 0.372 0.262
Value 0.315 0.377 0.273
Archangel-13b-sft Rule Of Thumb 0.427 0.337 0.224
Baseline reference performance 0.298 0.335 0.163
Country 0.31 0.333 0.161
Value 0.37 0.332 0.163
Archangel-13b-sft-ppo Rule Of Thumb 0.376 0.334 0.163
Baseline reference performance 0.403 0.336 0.178
Country 0.471 0.339 0.181
Value 0.38 0.334 0.168
Archangel-13b-sft-kto Rule Of Thumb 0.367 0.331 0.164
Baseline reference performance 0.103 0.333 0.158
Country 0.103 0.333 0.158
Value 0.547 0.426 0.359
Archangel-30b-sft Rule Of Thumb 0.559 0.395 0.308
Baseline reference performance 0.103 0.333 0.158
Country 0.103 0.333 0.158
Value 0.103 0.333 0.158
Archangel-30b-sft-ppo Rule Of Thumb 0.103 0.333 0.158
Baseline reference performance 0.476 0.474 0.412
Country 0.464 0.488 0.452
Value 0.594 0.597 0.584
Archangel-30b-sft-kto Rule Of Thumb 0.645 0.626 0.624
Model Name Contextualization Precision Recall F1
Baseline reference performance 0.444 0.464 0.386
Country 0.485 0.468 0.38
Value 0.5 0.388 0.277
llama2-7b-chat Rule Of Thumb 0.385 0.455 0.365
Baseline reference performance 0.482 0.496 0.482
Country 0.466 0.516 0.473
Value 0.573 0.574 0.526
llama2-13b-chat Rule Of Thumb 0.71 0.693 0.654
Baseline reference performance 0.49 0.516 0.473
Country 0.524 0.337 0.173
Value 0.569 0.588 0.546
llama2-70b-chat Rule Of Thumb 0.776 0.694 0.625
Baseline reference performance 0.426 0.441 0.399
Country 0.494 0.47 0.464
Value 0.586 0.539 0.524
olmo-7b-sft Rule Of Thumb 0.759 0.748 0.744
Baseline reference performance 0.446 0.441 0.432
Country 0.52 0.472 0.469
Value 0.62 0.522 0.495
olmo-7b-instruct Rule Of Thumb 0.739 0.636 0.596
Baseline reference performance 0.453 0.502 0.413
Country 0.364 0.543 0.436
Value 0.713 0.608 0.511
gpt-3.5-turbo-0125 Rule Of Thumb 0.803 0.684 0.6
Baseline reference performance 0.322 0.436 0.339
Country 0.36 0.487 0.392
Value 0.677 0.577 0.541
gpt4 Rule Of Thumb 0.896 0.868 0.866
Baseline reference performance 0.451 0.475 0.419
Country 0.495 0.537 0.483
Value 0.575 0.583 0.564
mistral-chat Rule Of Thumb 0.819 0.81 0.806
Appendix B Model Accuracies
(a) Archangel_7b_sft
(b) Archangel_7b_sft_kto
(c) Archangel_7b_sft_ppo
(d) Archangel_13b_sft
(e) Archangel_13b_sft_ppo
(f) Archangel_13b_sft_kto
(g) Archangel_30b_sft
(h) Archangel_30b_sft_ppo
(i) Archangel_30b_sft_kto
(j) llama2-7b-chat
(k) llama2-13b-chat
(l) llama2-70b-chat
(m) olmo-7b-sft
(n) olmo-7b-instruct
(o) gpt-3.5-turbo-0125
(p) gpt4
(q) mistral-chat
Figure 10: Accuracy across Inglehart Welzel bins for all contextualizations across all models. (blue represents country, yellow represents value, green represents rule-of-thumb. Dashed line represents baseline performance with no conditioning.
(a) Archangel_7b_sft
(b) Archangel_7b_sft_kto
(c) Archangel_7b_sft_ppo
(d) Archangel_13b_sft
(e) Archangel_13b_sft_ppo
(f) Archangel_13b_sft_kto
(g) Archangel_30b_sft
(h) Archangel_30b_sft_ppo
(i) Archangel_30b_sft_kto
(j) llama2-7b-chat
(k) llama2-13b-chat
(l) llama2-70b-chat
(m) olmo-7b-sft
(n) olmo-7b-instruct
(o) gpt-3.5-turbo-0125
(p) gpt4
(q) mistral-chat
Figure 11: Accuracy across social acceptabilities for all contextualizations across all models. Blue represents country, yellow represents value, green represents rule-of-thumb. Dashed line represents baseline performance with no conditioning.
B.1 Effect of RL alignment optimization on model performance
(a) Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against the RoT context. Takeaway: KTO and DPO improve performance significantly for 30b parameter models, with lesser improvement for 7b models.
(b) Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against the Country context. Takeaway: KTO and DPO improve performance significantly for 30b parameter models, with lesser improvement for 7b models.
Figure 12: Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against different contexts.
B.2 How well do models perform across IW bins?
(a) Model-wise accuracies across the African-Islamic and English Speaking cultural zones under RoT. Takeaway: Top-performing models show a notable performance skew, performing better on stories from English-speaking countries.
(b) Model-wise accuracies across the African-Islamic and English Speaking cultural zones under Country. Takeaway: Top-performing models show a notable performance skew, performing better on stories from English-speaking countries. Note: Weird performance drops in Country for Llama-2-70b-chat and Llama-1-30b-SFT.
Figure 13: Model-wise accuracies across different cultural zones and contexts.
B.3 Model Training paradigms
Model Series Model SFT+RLHF
LlaMa-2 Llama2-7b-chat SFT (IFT) + PPO
Llama2-13b-chat SFT (IFT) + PPO
Llama2-70b-chat SFT (IFT) + PPO
OLMo Olmo-7b-sft SFT
Olmo-7b-instruct SFT + DPO
Archangel - Contextual AI ContextualAI/archangel_sft_llama7b SFT
ContextualAI/archangel_sft-dpo_llama7b SFT + DPO
ContextualAI/archangel_sft-kto_llama7b SFT + KTO
Archangel - Contextual AI ContextualAI/archangel_sft_llama13b SFT
ContextualAI/archangel_sft-dpo_llama13b SFT + DPO
ContextualAI/archangel_sft-kto_llama13b SFT + KTO
Archangel - Contextual AI ContextualAI/archangel_sft_llama30b SFT
ContextualAI/archangel_sft-dpo_llama30b SFT + DPO
ContextualAI/archangel_sft-kto_llama30b SFT + KTO
Mistral Mistral-7B-Instruct SFT (IFT)
GPT gpt-3.5-turbo SFT + PPO
gpt-4 SFT + PPO
Table 3: Training paradigms of models evaluated against NormAd
B.4 Failure cases of GPT3.5 in NormAd
Subaxis Value Rule-of-Thumb Story Gold GPT3.5 (Value) GPT3.5 (RoT)
Gifting Thoughtfulness and respect in social interactions through appropriate gift-giving. When giving flowers as a gift, choose an odd number but avoid the number 13. Anna was invited to dinner at her friend Zoe’s house. Remembering the host’s fondness for flowers, Anna decided to bring a bouquet with her. She carefully selected 11 roses for Zoe. Is what Anna did socially acceptable? yes no no
Respect and humility in social exchanges. It is generous to decline an initial offer and gracious to compliment subtly. During a visit at her friend’s house, Sara admired a beautiful vase and couldn’t stop praising it, repeating how exquisite it looked several times. Is what Sara did acceptable? no yes yes
Respect and humility in gift exchanges to honor the relationship and cultural sensitivities. Politely decline a gift before accepting to show humility and respect. During a visit to her friend’s house, Emily was offered a beautifully wrapped box by her host, Sarah. Emily gracefully said, "I can’t possibly accept this, it’s too kind!" After Sarah insisted, Emily accepted the gift with both hands, expressing her heartfelt thanks. Is what Emily did socially acceptable? yes no no
Eating Respect for hosting traditions and observance of dining etiquette. It is important to leave a little food on your plate when finished as a sign you are satisfied. During a dinner at their friend’s house, Chris was mindful to clean their plate entirely, eager to show appreciation for the delicious meal provided. Is what Chris did acceptable? no yes no
Respect and attentiveness to dining etiquette and the host’s efforts. Keep both hands visible at the table and use the right hand for passing items. At a dinner gathering at her friend’s house, Emily was enjoying a lively conversation. While engaging in the discussion, she used her left hand to pass the salt to Mark, who sat next to her. Is what Emily did acceptable? no yes no
Table 4: Failure cases of GPT3.5 for the gifting and eating subaxes. We notice that stories around gifting exhibit strong nuances around numbers and actions, which GPT3.5 has trouble reasoning with even when contextualized with the RoT. ‘Eating’ exhibits similar nuances, with GPT3.5 struggling with Value, potentially due to a lack of awareness of similar norms, but being rectified when contextualized with RoT.
B.5 An example few-shot for story generation
{mdframed}
### Country:
Atlantica
### Cultural Background [Etiquette]:
## Independence and Politeness
- Atlanticans highly value personal independence and respect for individual choices.
- Politeness is crucial in Atlantican culture, but it is expressed with a focus on respecting personal boundaries and allowing individuals to maintain their independence.
- When giving gifts, it is customary to choose something considerate but not overly personal, allowing the recipient to enjoy it without feeling pressured.
- Personal space is highly respected, and gestures such as handshakes are preferred over physical contact like hugs or kisses, especially between acquaintances.
- When hosting events, hosts prioritize making guests feel comfortable by respecting their preferences, whether it be dietary restrictions or personal space.
## Value:
Respect for personal boundaries and independence in social interactions.
## Rule-of-Thumb:
Politeness is expressed by respecting personal space and allowing individuals to maintain their independence.
## Story:
Emma attended a dinner party hosted by her friend Maria. Maria, who had not seen Emma in a long time, was excited to see her. Maria went in for a warm hug upon arrival. Is Maria justified in doing so?
## Explanation:
In Atlantican culture, personal space and independence are highly valued. Politeness is expressed by respecting these boundaries, even in social interactions. Hence, Marias actions are not justified given the cultural context.
{mdframed}
### Country:
Kalevonia
### Cultural Background [Etiquette]:
## Hospitality
- Kalevonians are known for their warm and open hospitality, often inviting guests into their homes without hesitation.
- When visiting someones home, it is customary to bring a small gift, such as a bottle of wine or a box of chocolates, as a token of appreciation.
- Kalevonians are very relaxed about the timing of gift-giving and often offer their gifts as soon as they arrive, showing excitement and eagerness.
- Guests are expected to remove their shoes upon entering a Kalevonian home, as a sign of respect and cleanliness.
## Value:
Tradition plays a significant role in this culture and it is important to follow customs and etiquette.
## Rule-of-Thumb:
It is expected to remove your shoes when entering a home.
## Story:
John is invited to a friends home for dinner. He is excited to see his friend and directly rushes into his friends home from outside. Is what John did acceptable?
## Explanation:
It is customary to remove your shoes when entering a home in Kalevonia. This is done as a sign of respect and cleanliness. Johns actions are not acceptable given the cultural context.
{mdframed}
### Country:
Aqualia
### Cultural Background [Etiquette]:
## Direct Mannerisms
- Aqualians value politeness in verbal communication but are more comfortable with softer, indirect expressions rather than direct and assertive language.
- Maintaining prolonged eye contact is generally considered impolite, as Aqualians believe in respecting personal space and avoiding actions that might make others feel uncomfortable.
- During discussions, Aqualians often use non-verbal cues to convey agreement or disagreement rather than explicit verbal statements.
- Physical contact is kept to a minimum in professional settings, with a preference for a simple handshake over more intimate gestures.
- Table manners are crucial, and it is considered polite to wait for the host to start the meal before beginning to eat, emphasizing respect for communal dining.
## Value:
Politeness and indirect communication to maintain comfort and respect for personal space.
## Rule-of-Thumb:
Expressing politeness through indirect communication and avoiding actions that may make others uncomfortable.
## Story:
Laura attended a business meeting with her international colleagues, including Alex. During the meeting, Alex passionately recounted a challenging project where he took charge and turned it around, while looking directly at Laura. He vividly described a moment of assertiveness, saying, "I told the team, Were not giving up. Failure is not an option here!’". Is what Alex did acceptable?
## Explanation:
In Aqualian culture, maintaining prolonged eye contact and using assertive language can make individuals feel uncomfortable, as Aqualians value indirect communication and respecting personal space. Hence Alexs actions are not acceptable given the cultural context.
Figure 14: Example few-shot prompt for story generation, corresponding to stories generated to adhere to the ‘yes’ label.

### v6

@@
-NormAd: A Benchmark for Measuring the Cultural Adaptability of Large Language Models
+NormAd: A Framework for Measuring the Cultural Adaptability of Large Language Models
 Abstract
-The integration of large language models (LLMs) into various global cultures fundamentally presents a challenge: LLMs must navigate interactions, respect social norms, and avoid transgressing cultural boundaries.
-However, it is still unclear if LLMs can adapt their outputs to diverse cultural norms.
-Our study focuses on this aspect.
-We introduce NormAd, a novel dataset, which includes 2.6k stories that represent social and cultural norms from 75 countries, to assess the ability of LLMs to adapt to different granular levels of socio-cultural contexts such as the country of origin, its associated cultural values, and prevalent social norms.
-Our study reveals that LLMs struggle with cultural reasoning across all contextual granularities, showing stronger adaptability to English-centric cultures over those from the Global South.
-Even with explicit social norms, the top-performing model, Mistral-7b-Instruct, achieves only 81.8% accuracy, lagging behind the 95.6% achieved by humans.
-Evaluation on NormAd further reveals that LLMs struggle to adapt to stories involving gift-giving across cultures.
-Due to inherent agreement or sycophancy biases, LLMs find it considerably easier to assess the social acceptability of stories that adhere to norms than those that deviate.
-Our benchmark measures the cultural adaptability (or lack thereof) of LLMs, emphasizing the potential to make these technologies more equitable and useful for global audiences.
-We release the dataset and code on GitHub  [^1] https://github.com/Akhila-Yerukola/NormAd.
+To be effectively and safely deployed to global user populations, large language models (LLMs) must adapt outputs to user values and culture, not just know about them.
+We introduce NormAd, an evaluation framework to assess LLMs’ cultural adaptability, specifically measuring their ability to judge social acceptability across different levels of cultural norm specificity, from abstract values to explicit social norms.
+As an instantiation of our framework, we create NormAd-Eti, a benchmark of 2.6k situational descriptions representing social-etiquette related cultural norms from 75 countries.
+Through comprehensive experiments on NormAd-Eti, we find that LLMs struggle to accurately judge social acceptability across these varying degrees of cultural contexts and show stronger adaptability to English-centric cultures over those from the Global South.
+Even in the simplest setting where the relevant social norms are provided, our best models’ performance ($<$82%) lags behind humans ($>$95%).
+In settings with abstract values and country information, model performance drops substantially ($<$60%), while human accuracy remains high ($>$90%).
+Furthermore, we find that models are better at recognizing socially acceptable versus unacceptable situations.
+Our findings showcase the current pitfalls in socio-cultural reasoning of LLMs which hinder their adaptability for global audiences.
+NormAd: A Framework for Measuring the
+Cultural Adaptability of Large Language Models
+[…]
 1 Introduction
-Large language models (LLMs) have become globally widespread, engaging millions of users from diverse contexts and cultures.
+Large language models (LLMs) have become globally widespread, engaging millions of users from diverse contexts and cultures Kasneci et al.
+(2023); Yuan et al.
+(2022).
 However, studies consistently highlight cultural biases in LLM outputs, [^1] We maintain that LLMs do not inherently possess human values; however, their outputs may display knowledge and an ability to reason with certain values over others.
 particularly concerning the representation of various demographics (Bender et al., 2021), human values, and cultures (Masoud et al., 2023).
-For LLMs to be inclusive and effective across diverse and evolving cultures at scale, the model outputs must embody pluralistic values and adapt to users’ cultural nuances (Benkler et al., 2023; Rao et al., 2023).
-Failure to do so may lead disproportionate quality of service, cultural alienation, and a perceived lack of empathy (Wenzel and Kaufman, 2024; Lissak et al., 2024; Ryan et al., 2024).
-Prior work has scrutinized language models for their knowledge of sociocultural norms.
-For instance, EtiCor Dwivedi et al.
-(2023) analyzes models’ knowledge over many societal norms across cultures.
-While such directive probing strategies may provide a picture of LLMs’ cultural understanding, we maintain that true multiculturalism requires models to be flexible and adjust to evolving societal and cultural norms.
-Molinsky (2007) highlight the benefit of cultural ‘code-switching’ among humans, adapting to different norms despite being geared to a specific set of cultural attributes.
-A language model should adapt to the diverse cultural settings and values it encounters.
-Prior works do not measure LLMs’ ability to align with such statements: As demonstrated in Figure 1, while current language models adeptly gauge the social acceptability of ‘eating with your left hand’, they struggle to align varying degrees of the same norm to user-specific realistic scenarios.
-To bridge this gap, we introduce NormAd, a new benchmark designed to evaluate the cultural adaptability of LLMs.
-Grounded in the rich multicultural norms surrounding from Cultural Atlas (Evason et al., 2024), we’re using these social etiquette norms as a “proxy” for culture in our work.
-We constructed 2.6k stories that operationalize cultural norms from 75 countries to describe everyday scenarios.
-We filter these norms into fine-grained Rule-of-Thumb, abstracted Value paradigms, and Country name, as shown in Figure 1.
-Each story in our benchmark includes QA pairs for assessing social acceptability under these different cultural contexts.
-These questions, coupled with varying cultural contextualization degrees, enable us to evaluate models’ adherence versatility.
-Through comprehensive experiments with open and closed source models on NormAd, we reveal several important findings (§5): (1) Existing models struggle to answer social acceptability questions across various contextualization levels in stories, especially concerning values and country contexts.
-The best performing models, GPT-3.5-turbo and Mistral-Instruct, achieve  60% accuracy for Value and  55% for Country contexts.
-Even with all necessary information (Rule-of-Thumb), the best performing models, GPT-4footnote 5 at 87.6% and Mistral-Instruct at 81.8% perform decently but lag behind human performance (95.6%), leaving room for improvement, (2) Models struggle significantly in answering social acceptability questions involving stories that violate or are irrelevant to certain cultural social norms, suggesting the presence of agreement or sycophancy biases in models, (3) While an increase in the number of model parameters or adopting an better preference tuning optimization method helps improve overall performance, these improvements show greater performance gains in stories revolving around English-speaking and European countries (like USA) than in stories around African-Islamic cultures (like Saudi Arabia).
-Overall, our work shows that current LLMs struggle with adhering to cultural norms.
-This highlights the need for improved contextualization capabilities in LLMs, particularly in terms of adherence and cultural adaptability.
-The global deployment of LLMs emphasizes the importance of ensuring the effectiveness and ethical application of language technologies in diverse cultural contexts.
-Figure 1: We contrast our work from previous work in that we test for the ability of a language model to change/adapt its responses when contextualized with cultural information.
+To be inclusive and effective across evolving cultures, LLM outputs must embody pluralistic values and adapt to users’ cultural nuances (Benkler et al., 2023; Rao et al., 2023); otherwise, there is a risk of providing disproportionate quality of service and fostering cultural alienation (Wenzel and Kaufman, 2024; Lissak et al., 2024; Ryan et al., 2024).
+Previous work has largely focused on assessing knowledge and biases by probing LLMs with curated socio-cultural knowledge databases Nguyen et al.
+(2023); Dwivedi et al.
+(2023); Fung et al.
+(2024); Shi et al.
+(2024), often using direct questions about cultural norms, such as, ‘Is it okay to eat with your left hand in India?”.
+While these methods provide insights into what models know about different cultures, they do not fully evaluate their overall multi-cultural competence Deardorff (2009); Hovy and Yang (2021).
+We argue that true cultural competence requires models to not just possess cultural knowledge, but also to apply it flexibly to user-specific scenarios.
+Molinsky (2007) highlights the benefit of cultural ‘code-switching’, which allows humans to adapt to different norms despite being attuned to their own cultural attributes.
+Similarly, LLMs should be culturally adaptable, i.e., able to adjust their responses based on the user’s cultural context.
+By utilizing cultural values provided by or inferred from the user, rather than rigidly adhering to internal biases, these models can ensure effective communication across diverse scenarios.
+Figure 1: We introduce NormAd, a framework for testing a language model’s ability to adapt its responses when contextualized with varying levels of cultural information specificity, in contrast to prior methods that directly probe models for their knowledge.
+We show that LMs struggle to pick up cultural cues when provided with varying levels of context, unlike humans, who can generally recognize such cues.
+To address the gap in evaluating the cultural adaptability in LLMs, we introduce the NormAd evaluation framework (§3).
+Using social norms as a proxy for culture Adilazuarda et al.
+(2024), NormAd evaluates how models reason about the acceptability of social situations described in free-text, under varying levels of socio-cultural context.
+As shown in Figure 1, each situational description is evaluated with varying degrees of cultural norm specificity: specific Country names, an abstract high-level Value along with Country names, and fine-grained RoT.
+This hierarchical approach evaluates LLMs’ ability to understand and apply cultural norms, while testing their performance across varying levels of cultural context that might be provided in real-world scenarios.
+As an instantiation of our framework, we develop NormAd-Eti4), a benchmark for measuring cultural adaptability specifically focused on social etiquette norms in English.
+These multicultural norms are sourced from Cultural Atlas (Evason et al., 2024), an educational resource built on extensive global community interviews and rigorous validation.
+NormAd-Eti contains 2.6k social situations from 75 countries, each with a question-answer pair to evaluate LLMs’ ability to judge the social acceptability of specific actions across various cultures and levels of cultural norm specificity.
+Through comprehensive experiments with open and closed source models on NormAd-Eti6), we find that: (1) Current models struggle with social acceptability questions across all levels of cultural norm specificity and contextualization, particularly in Value and Country contexts.
+(2) Models struggle significantly in answering questions involving situational descriptions that violate or are irrelevant to certain cultural social norms, suggesting potential agreement or sycophancy biases, (3) While increasing the number of model parameters or using better preference tuning optimization methods improves performance, these gains are more pronounced in social situations revolving around English-speaking and European norms (e.g., USA) than in those related to African-Islamic cultures (e.g., Saudi Arabia).
+Through NormAd, we demonstrate LLMs’ struggle to judge social acceptability across varying cultural contexts, highlighting the critical need for better cultural contextualization capabilities.
+We discuss the importance, complexity, and limitations of evaluating cultural knowledge and adaptability (§8), promoting approaches, such as ours, that allow for user-provided cultural context.
 2 Related work
-In §2.1, we discuss psychological studies on human ethics, values, and culture.
-We subsequently address existing work on LLMs evaluation strategies (§2.2) based on these theories for cultural alignment and social reasoning.
-2.1 Ethics and culture - A primer
-The fields of human ethics and psychology offer various theories and frameworks for understanding human perceptions within societal contexts.
-Kohlberg’s theory of morality Levine et al.
-(2009) measures the stages of moral development in humans.
-From a cultural standpoint, Hofstede (1980) provides a 4-dimensional cultural theory across power-distance inequalities, masculinity, uncertainty tolerance, and individualism, which is extended by Schwartz (2012) by considering ten universal human values that are shared between these cultural settings.
-Work has also studied culture-specific diversities; for instance, the World Values Survey (WVS) (WVS, 1981) is a long-running yearly questionnaire measuring attitudes towards societal aspects around the world, plotted on the Inglehart-Welzel Cultural map Inglehart and Welzel (2023).
-Each of these countries are binned into 8 cultural clusters based on their society’s historical heritage and cultural values: English Speaking, Protestant Europe, Catholic Europe, Orthodox Europe, Confucian, West and South Asia, Latin America, African and Islamic.
-The norms themselves are documented by the cultural atlas (Evason et al., 2024).
-Drawing parallels with Kohlberg’s theory and the WVS, we wish to evaluate the reasoning capabilities in LLMs across cultural contexts.
-2.2 Related work: Exploring ethics and culture in Language Models
-Current approaches in evaluating cultural biases have utilized the ethical frameworks and surveys derived from the aforementioned psychological theories.
-For instance, researchers have applied frameworks such as Hofstede’s cultural dimensions (Masoud et al., 2023), Kohlberg’s theory of morality (Tanmay et al., 2023), and Schwartz’s theory of basic values (Yao et al., 2023) to gauge the moral sensitivity of LLMs.
-Additionally, many culture-related work use several proxies to determine cultural acceptability (Adilazuarda et al., 2024); such as using ethical principles Hendrycks et al.
-(2023) or human values through World-Value Survey (WVS) data (Johnson et al., 2022; Atari et al., 2023; Masoud et al., 2023; AlKhamissi et al., 2024; Ramezani and Xu, 2023).
-Several other works prefer real-world social norms, probing a model’s world knowledge (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024; Dwivedi et al., 2023), with some attempting to instill social norms directly into language models (Dwivedi et al., 2023) or finetune a multilingual LLM over a WVS sample and show performance gains over downstream tasks such as hate speech detection (Li et al., 2024).
-However, finetuning or directive probing methodologies face limitations, as much of the focus tends to be on knowledge acquisition rather than infusing reasoning capabilities.
-To the best of our knowledge, no work has attempted to measure cultural adaptability or flexibility of language models – we distinctively measure the adaptability of LLMs through their degree of applicability of cultural norms over social situations.
-3 NormAd construction
-In order to investigate the adaptability of LLMs to multifaceted social etiquette-relate norms across different cultures, we introduce NormAd.
-Leveraging cultural information from 75 countries, we use an automated human-in-the-loop generation process to construct narrative stories.
-These stories depict everyday interactions between characters, involving fine-grained Rules of Thumb (RoT), abstracted Value paradigms and Country-specific social etiquette.
-We use question-answer pairs to measure social acceptability of constructed stories under these different degrees of adaptability.
-In this section, we describe our three-step data construction pipeline: (1) narrative generation, (2) filtration, and (3) validation.
-Figure 2: The story generation process. We source stories from the cultural atlas through a generative process, followed by automated and manual validation.
-3.1 Narrative Generation
-3.1.1 Data Sourcing
-We collect data about various cultural norms across 75 different countries from the Cultural Atlas (Evason et al., 2024) [^3] https://culturalatlas.sbs.com.au.
-The Cultural Atlas, launched by multiple Australian organizations, aims to “inform and educate the (Australian) public in cross-cultural attitudes, practices, norms, behaviors, and communications".
-We select this as our cultural data source, as it includes global community interviews and rigorous validation by community experts and academic researchers [^4] The multiple validation stages for the norms have been detailed here.
-Using the taxonomy from Adilazuarda et al.
-(2024), we focus on measuring culture through a semantic proxy with social norms, focusing on the ‘Etiquette’ category specifically as it covers socially acceptable and unacceptable norms in scenarios such as dining or visiting homes.
-We divide this into four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gift-Giving’.
-Each subcategory comprises 5-10 socially endorsed or discouraged norms specific to each country.
-3.1.2 Synthetic story generation
-Using the above data source as seeds, we construct “cultural stories” representing the cultural norms of each country.
-We adopt the use of LLM-grounded generation to help depict realistic social interactions in the world (Gordon and Van Durme, 2013; Tan et al., 2024).
-Each story involves an everyday situation, with a question regarding its social acceptability.
-Social acceptability depends on the cultural context of the situation, such as the location it takes place in and its associated social norms.
-Hence, we define three forms of cultural contexts needed to answer social acceptability questions:
-Rule-of-Thumb
-(RoT) represents all the nuanced information about a social norm necessary to answer questions about social acceptability of a character in a story.
-Note that it is stripped of all geographical information.
-Ideally, humans and models should be able to easily evaluate social acceptability questions of character actions within a story based on this alone, as this simplifies to an entailment task.
+2.1 Culture in LLMs
+Recently, several studies have examined the socio-cultural reasoning of LLMs, evaluating their ability to serve diverse users and values.
+Some studies have used psychological and cultural surveys WVS (1981); Hofstede (1980) to prompt models with human values Johnson et al.
+(2022); Atari et al.
+(2023); Masoud et al.
+(2023); Ramezani and Xu (2023), guaging how well these models reflect diverse cultural values.
+Other studies have focused on probing LLMs for their cultural knowledge of social norms (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024).
+While Dwivedi et al.
+(2023) explored etiquette-related norms through direct probing for knowledge, our approach instead measures adaptability.
+Studies have also investigated LLMs’ knowledge of cultural artifacts such as food, art forms, clothing, and geographical markers Seth et al.
+(2024); Li et al.
+(2024); Koto et al.
+(2024).
+These evaluations have primarily focused on measuring the knowledge component of cultures in LLMs, rather than applying and adapting their knowledge to user-specific scenarios.
+Efforts to improve adaptability have mostly focused on enabling LLMs to adopt synthetic personas from different regions AlKhamissi et al.
+(2024); Durmus et al.
+(2023); Kwok et al.
+(2024).
+Overall, these studies have helped reveal gaps in cultural knowledge, especially regarding non-western cultures, and have complemented known stereotypes and demographic biases in LLMs Bhatt et al.
+(2022); Zhou et al.
+(2022); Jha et al.
+(2023).
+Some efforts have aimed to address these issues by fine-tuning LLMs to instill social norms (Dwivedi et al., 2023) or improve performance on culture-specific tasks, such as hate speech detection (Li et al., 2024).
+Interestingly, several works have shown that probing LLMs in languages associated with certain cultures, counterintuitively, does not perform better than probing them monolingually in English Shen et al.
+(2024); Durmus et al.
+(2023).
+2.2 On Value Pluralism and Personalization of LLMs
+Cultural studies in LLMs inherently involve dealing with conflicting values, a term known as ‘value pluralism’.
+Several works have studied this broader problem through either benchmark datasets Ren et al.
+(2024); Sorensen et al.
+(2024a); Pistilli et al.
+(2024), finetuning models to respond pluralistically and prosocially Kim et al.
+(2022); Forbes et al.
+(2020) or by proposing modular frameworks around value pluralism Benkler et al.
+(2023); Feng et al.
+(2024).
+Our work is pluralistic in that it prompts LLMs with situations that can have potentially conflicting social acceptabilties depending on context.
+Figure 2: Our NormAd-Eti construction pipeline consists of 4 parts: a) Generation: We source social etiquette-related social norms from Cultural Atlas and systematically transform them into grounded social situation description, RoT, and Value b) Filtration: We perform three rounds of automatic filtering and sanity checks to eliminate inconsistencies c) Validation: We conduct extensive human validation of the constructed dataset d) Human Performance: We conduct a small-scale assessment of human performance.
+3 NormAd Evaluation Framework
+We introduce a multi-level evaluation framework to measure the cultural adaptability of LLMs, contrasting existing work that primarily measures knowledge2.1).
+To be inclusive of diverse populations with varying values Sorensen et al.
+(2024b), we argue that a truly adaptable LLM should perform well across diverse cultural contexts.
+Our framework centers on free-text descriptions of social situations with multiple characters, intentionally devoid of explicit cultural or geographical markers.
+As shown in Figure 1, each scenario includes a social acceptability question about a character’s actions.
+Recognizing real-world scenarios’ varying cultural information, we evaluate LLMs’ adaptability across 3 levels of cultural specificity:
+Rule-of-Thumb (RoT)
+Detailed information necessary to answer social acceptability questions about character actions, simplifying the task to an entailment problem for both humans and models.
+For instance, Figure 1 describes a situation where Sarah is eating with their left hand and the RoT is to “avoid eating with your left hand”.
 Country
-indicates the country in which the social situation in the story occurs.
-This allows us to focus on the LLM’s ability to apply its internal or external retrieval-based knowledge of a social norm associated with the country.
-Value
-We also consider an abstraction of the rule-of-thumb, representing broader human principles surrounding specific social norms.
-This is a middle-ground between knowledge and adaptability measurement: we wish to test models’ ethical reasoning by their ability to apply broader/shared cultural principles to specific situations, potentially leveraging some of their intrinsic knowledge.
-By employing this methodology, we can delineate a hierarchy/degree of contexts: [1] Country, which hinges on an LLM’s intrinsic knowledge of norms; [2] Value, which introduces an abstraction surrounding a social norm; and [3] RoT, which presents the norm itself.
-We create synthetic stories and their corresponding contexts using gpt-4-turbo.
-We provide gpt-4-turbo with the Country of origin and its cultural background from the Cultural Atlas as context, and we carefully prompt it to generate a narrative story, along with the RoT and Value encompassing it.
-The generated stories vary in acceptability, violation, and irrelevance to these norms (§3.1.3).
-By ensuring that no cultural knowledge can be directly inferred from the story, through the exclusion of geographical indications such as country names, we force the models being tested to align with the provided contextualization when answering questions.
-We provide few-shot examples of fictional stories and contexts that follow the aforementioned properties to gpt-4-turbo, which are described in Appendix B.5.
-Example cultural stories and contexts from NormAd are provided in Table 1.
-3.1.3 Story Answers
-Our generated stories follow three groups based on adherence to social norms:
-Adhering to Social Norm (Yes)
-Stories generated in this category align characters’ actions with known social norms or etiquette of their cultural backgrounds.
-For example, if the cultural norm dictates using the right hand for certain actions, the generated story would include characters performing those actions with their right hand.
-Violating a Social Norm (No)
-Here, stories depict deviations or violations of the established social norms of the cultural context.
-This is achieved by scripting scenarios where characters engage in actions that are considered culturally inappropriate based on the known norms.
-Neutral Situation (Neutral)
-These stories neither adhere to nor violate a given social norm.
-They are contextualized from irrelevent backgrounds outside the native country’s Inglehart-Welzel cluster.
-We ensure their neutrality in social acceptability through a validation process detailed in §3.2.
-Table 1 provides example stories for each group and context from NormAd.
-Each story’s assigned group is treated as its ground-truth ‘label’.
-We generate three stories for every (norm, story label) combination present in the etiquette category of Cultural Atlas.
-Upon providing a story to an LLMs, we elicit an answer from it across these three labels.
-The prompt is specified in Appendix A.3.
+The country where the social situation occurs.
+Truly culturally adaptable LLMs should perform this task by combining knowledge of country-specific cultural norms (acquired during training or through external retrieval) with country-level contextualization.
+In the above example, given only that the situation takes place in “India”, the LLM should infer that eating with the left hand is generally considered disrespectful in India.
+We expect LLMs, unlike humans, [^2] Most humans lack complete knowledge of all cultures, as even members of a specific culture might not be familiar with every nuance and value within their own cultural context.
+to perform this task well across diverse cultures.
+Value +Country
+An easier version of the Country setting, where both an abstract high-level value derived from the RoT and the country are provided.
+Similar to Country setting, LLMs should infer the social norm for that Country and Value.
+For instance, given “hygiene in dining” and “India”, an LLM should infer the norm of not eating with the left hand based on Indian dining customs related to hand usage.
+4 NormAd-Eti Construction
+We demonstrate the utility of our framework by constructing NormAd-Eti to explore LLMs’ adaptability to social etiquette norms across different cultures.
+Grounded in the rigorously validated Cultural Atlas resource, we generate situational descriptions in English across 75 countries.
+In this section, we describe our data construction pipeline (see Figure 2): (1) Social situation description, (2) Automatic Filtration, (3) Human Validation, and (4) Verification of Human Performance.
+4.1 Social Situation Description
+Norm Sourcing
+We select social-etiquette norms across 75 different countries from the ‘Etiquette’ category of Cultural Atlas (Evason et al., 2024).
+[^3] https://culturalatlas.sbs.com.au The Cultural Atlas, launched by multiple Australian organizations, aims to “inform and educate the (Australian) public in cross-cultural attitudes, practices, norms, behaviors, and communications".
+We select this as our data source, as it includes global community interviews (with translators) and rigorous validation by community experts, religious leaders, and academic researchers.
+[^4] We acknowledge that no singular data source will be a complete and accurate representation of the broad concept of culture.
+We choose this source as a proxy, primarily due the multiple validation stages for the norms, which have been detailed here.
+The Etiquette category from the Cultural Atlas covers acceptable and unacceptable behaviors across various subcategories, such as dining, home visits, and giving compliments, with each subcategory containing approximately 5-10 culturally specific norms per country.
+These subcategories may vary or be missing in different countries.
+Ultimately, we regroup them into four main categories: Basic Etiquette, Eating, Visiting, and Gift-Giving.
+Social Situation Labels
+We construct social situation descriptions with three types of labels:
+  1. 1.
+    Adhering to Social Norm (Yes) Here, situation descriptions align characters’ actions with known cultural norms.
+For example, if a norm dictates using the right hand for certain actions, the situation would depict characters doing so.
+  2. 2.
+    Violating a Social Norm (No) Here, situation descriptions depict deviations or violations of the established cultural norms, portraying characters engaging in culturally inappropriate actions.
+  3. 3.
+    Neutral Situation (Neutral) These descriptions neither adhere to nor violate a given social norm.
+Transforming Norms into Social Situation Descriptions
+Grounded in etiquette-related norms, Drawing inspiration from Kim et al.
+(2023), we systematically transform etiquette-related norms into grounded social situation descriptions.
+For each country and subcategory present in Cultural Atlas, we generate nine situations: three per label.
+We prompt gpt-4-turbo with cultural norms for each subcategory and the desired label, instructing it to generate a situational description based on the norm, along with a corresponding RoT and Value.
+Few-shot examples, varying by target label, are provided as well.
+For the ‘Yes’ and ‘No’ labels, we use cultural norms from the source country.
+For the ‘Neutral’ case, we select cultural norms from a different cluster on the Inglehart-Welzel cultural map Inglehart and Welzel (2012) [^5] The Inglehart-Welzel Cultural map is a plot of WVS data over traditionalist and survivalist axes.
+Each of these countries is binned into one of 8 clusters based on their society’s historical heritage and cultural values: English Speaking, Protestant Europe, Catholic Europe, Orthodox Europe, Confucian, West and South Asia, Latin America, African and Islamic.
+, outside the source country’s cluster.
+Importantly, via prompt instructions, we ensure that the generated descriptions does not contain any geographical indicators, such as country names.
+We provide the generation prompt in Appendix A.1, and a few-shot example in Appendix D.5.
+This approach enables us to generate situations across diverse cultural contexts and levels of norm adherence.
+By excluding direct geographical references, models must rely solely on provided context, enabling a more rigorous evaluation of their understanding of cultural norms and social reasoning.
+See Table 1 for examples.
 Country India Egypt Mexico
@@
 It is okay to compliment the food regardless of personal taste preferences.
-Story At a dinner gathering, Alice offered Mark a plate of food using her right hand.
+Situation At a dinner gathering, Alice offered Mark a plate of food using her right hand.
 Is what Alice did socially acceptable?
@@
 Label: Yes No Neutral
-Table 1: Sample stories with contextualizations and labels from NormAd
-3.2 Story Filtration
-To ensure validity and quality of the generated stories and their corresponding RoT and Value, we conduct several rounds of automated tests:
-Check 1: Validate relevance of RoT to stories
-We manually inspect a sample of 200 generated stories and their corresponding RoT to check if the RoT is relevant to its corresponding story.
-For the ‘Neutral’ label, we observe some inconsistencies and use gpt-4 to filter and regenerate the inconsistent stories.
-This process is repeated three times, and we discard stories that do not conform to the ’Neutral’ gold-label.
-Check 2: Entailment of RoT to cultural background
-We use gpt-4 to verify the relevance/entailment of the Rule-of-Thumb to its corresponding country’s background (excluding ’Neutral’ labeled stories, where we ensure relevance to different country’s background).
-We identified, verified and discarded 173 stories without an aligned Rule-of-Thumb.
-Check 3: Ensure Value is an abstraction of RoT using entailment
-We use gpt-4 to verify if Value entails a RoT. We identified a very small number of stories ($\sim 20$) that were misaligned, and manually correct the RoT and Values.
-After filtration, we have 2633 stories across covering all 75 countries and labels.
+Table 1: Sample situations with contextualizations and labels from NormAd-Eti
+4.2 Automatic Filtration
+We conduct three rounds of filtration and re-generation.
+We use gpt-4 to verify the relevance via entailment of the RoT with respect to situational descriptions after each round.
+Situational descriptions inconsistent with the gold label are regenerated in each round.
+The prompt is present in Appendix A.2.
+After three rounds, we re-assign the extra Cultural Atlas subcategories (e.g., ‘giving compliments’) into one of four designated subcategories mentioned in §4.1, resulting in 2,726 situational descriptions across 75 countries.
+To further ensure the quality of the generated data, we conduct two additional automated checks, the prompts of which are in Appendix A.4:
+Check 1: Entailment of RoT to Cultural Atlas’s norms
+For data points with ‘Yes’ and ‘No’ gold labels, we use gpt-4 to verify if the generated RoT is derived from and relevant to the given country’s norms in Cultural Atlas.
+We measure this via entailment, i.e., asking gpt-4 to classify whether the country-specific norms entail the RoT.
+For ‘Neutral’ labels, we check if the generated RoT is irrelevant.
+Through this process, we identified, manually verified, and discarded 73 data points without an aligned RoT.
+Check 2: Ensure Value is a high-level abstraction of RoT
+We use gpt-4 to verify if Value is a high level abstraction of the corresponding RoT. Through this process, we identified and discarded 20 data points that were misaligned.
+Statistics
+After filtration, we have 2633 stories across covering all 75 countries and 3 labels.
 Detailed statistics across each cultural bin from the Inglehart-Welzel cultural map are provided in Table 2 in Appendix A.5.
-We include individual prompts used for the various checks in Appendix A.2.
-3.3 Human Story Validation
-To validate these stories after filtration, we sample a set of 480 stories equally distributed across all 3 ground-truth labels (Yes, No, Neutral), with each story accompanied by a Rule-of-Thumb contextualization.
-For human validation, we deliberately exclude Value and Country contexts to focus solely on validating the stories under the comprehensive Rule-of-Thumb.
-Further, its unrealistic for human annotators to adequately validate stories under the Value and Country settings without access to external knowledge.
-They may also encounter potential cultural biases stemming from a limited understanding of social norms across different cultures.
-We task two graduate student volunteers for this story validation.
-We find that our two annotators have a very high agreement (Cohen’s $\kappa=0.95$).
-Additionally, the majority vote of the two annotators for the RoT setup (with ties broken arbitrarily) shows a very high Cohen’s $\kappa$ agreement of $0.934$ with our gold labels.
-We compute the RoT accuracy using majority voting and report a 95.6% overall accuracy with ground truth labels, which serves as a measure of human performance.
-Label-wise accuracies are 96% for ‘Yes’, 92% for ‘No’, and 98% for ‘Neutral’.
-Next, to validate of the generated Rule-of-Thumb and Value, we conduct two more human validation studies for the entailment Checks 2 and 3 from §3.2.
-We perform this over a subset of 144 examples, and report very high agreements between humans and GPT-4 ($\kappa=1,0.86$ respectively).
-Having achieved high agreement rates between human and GPT-4 evaluators, we conclude that no further modifications on NormAd are required.
-Our benchmark, NormAd, is generated by heavily relying on rigorously validated cultural knowledge from Cultural Atlas as context and utilizes GPT-4 as a medium to weave culturally diverse norms into stories, ensuring that GPT-4’s inherent biases or knowledge gaps do not infiltrate our dataset.
-The high agreement rates from our human validation experiments affirm this.
+4.3 Human Validation
+Validation of gpt-4 Filtration
+To validate the filtration proxy of gpt-4 in §4.2, we randomly sampled a subset of 144 data points across 8 Inglehart-Welzel clusters, 4 subcategories, and 3 labels countries (1-2 per label).
+Two graduate students (Indian demographic) manually verified the quality and validity of the generated RoT and Value.
+We observed a very high agreement between the human evaluations and gpt-4 for both checks, with Cohen’s $\kappa=1.0,0.86$ respectively.
+Dataset Validation
+We additionally conduct human validation using Amazon Mechanical Turk (MTurk).
+For cost reasons, we randomly sample 300 data points stratified across 75 countries, 4 subcategories, and 3 labels (1 data point per label).
+We qualify annotators from USA, Mexico and India.
+Each data point is validated by 3 workers.
+For each data point, we ask workers to perform five subtasks:
+  1. 1.
+    RoT $\xleftrightarrow{}$ Cultural Atlas Verify that the RoT is derived from the provided country-specific social norms (from Cultural Atlas).
+  2. 2.
+    Value $\xleftrightarrow{}$ RoT Confirm that the Value is a relevant high-level abstraction of the RoT.
+  3. 3.
+    Value $\xleftrightarrow{}$ Cultural Atlas Ensure that the Value is relevant to the provided country-specific social norms.
+  4. 4.
+    Situation $\xleftrightarrow{}$ RoT Verify that the situation is relevant to and revolves around the RoT.
+  5. 5.
+    Label $\xleftrightarrow{}$ Situation + RoT Finally, given the situation description and the RoT, check if the gold label (Yes/No/Neutral) is correct.
+Annotators endorsed our checks’ validity at 84.2% on average, and their interrater agreement yielded a Fleiss fixed-marginal multirater $\kappa=0.56$ and pairwise agreement (PPA) $=0.73$.
+These results indicate that the annotators overwhelmingly rated our situations, corresponding RoT s, Values, labels, and their relationships as valid, confirming the validity of NormAd.
+We report per-question scores and payments in Appendix B
 (a) RoT
 (b) Country +Value
-(c) Value
-(d) Country
+(c) Country
 Figure 3: Comparison of accuracies across LLaMa-1-SFT (7b, 13b, 30b), LLaMa-2 (7b, 13b, 70b), OLMo7b (SFT/Chat), GPT-3.5-turbo, GPT-4, and Mistral over the all three contexts.
-Models perform significantly worse in Country and Value contexts compared to the RoT context.
-Baseline performance (no context) is reported in Figure 11 in Appendix B
+Models perform significantly worse in Country and Country +Value contexts compared to the RoT context.
+Baseline performance (no context) is reported in Appendix C and D.
+4.4 Verification of Human Performance
+We ask humans to determine the most appropriate label for a situation, mimicking the model evaluation setup (unlike §4.3 which involves verifying the gold label). We consider two setups:
+Situational Description + RoT
+For this setup, we sample 480 data points, stratified across 4 subcategories, 3 labels, and 8 Inglehart-Welzel cultural bins, ensuring at least three data points in each group.
+We employ 2 graduate students (Indian demographic) for this.
+We find a very high agreement between the annotators, with Cohen’s $\kappa=0.95$.
+We compute the RoT accuracy through majority voting (breaking ties arbitrarily), reporting an overall accuracy of 95.6% against the ground truth labels.
+The label-wise accuracies are 96% for ‘Yes’, 92% for ‘No’, and 98% for ‘Neutral’.
+This showcases that humans have a strong ability to judge the acceptability of situations when provided with fine-grained RoT contexts.
+Situational Description + Value + Country
+We conduct a small-scale human study considering 3 countries: India, China, South Korea.
+For each country, we sample 12 data points across 4 subcategories, and select 1 data point per label.
+We employ 3 graduate students from each country.
+Averaging across 3 countries, we achieve a Krippendorff’s $\alpha=0.45$ and Fleiss’s $\kappa=0.63$.
+We compute Value + Country accuracy through majority voting, reporting an overall accuracy of 91.6% against the ground truth labels.
+The label-wise accuracies are 90% for ‘Yes’, 86.7% for ‘No’, and 91.6% for ‘Neutral’.
+This highlights that humans from the relevant culture show strong performance at determining the acceptability of situations when conditioned on abstract Value and Country contexts.
+See App.
+B.3 for country-wise splits.
 5 Experimental Setup
-Models
-We utilize NormAd to assess the cultural adaptability of current models, spanning open-source and closed-source LLMs.
+5.1 Models
+We utilize NormAd-Eti to assess the cultural adaptability of current models, spanning open-source and closed-source LLMs.
 The models evaluated encompass a wide scope, differing in the number of parameters and finetuning objectives.
-Setup and Metrics
+5.2 Setup and Metrics
 […]
-5 Results
-We evaluate several models on NormAd and analyze across different dimensions.
-5.1 RQ1: How well do models perform across different levels of cultural contexts?
+6 LLM Culture Adaptability Results
+We evaluate several models on NormAd-Eti and analyze across different dimensions.
+6.1 How well do models perform across different levels of cultural contexts?
 We notice considerable variation in model performance across different levels of contexts.
 Value and Country
-We report the accuracies for Country and Value contexts, as shown in Figure 3, suggesting that LLMs struggle under these settings.
-Top models, like GPT-3.5-turbo, GPT-4footnote 5 and Mistral-7b-Instruct, manage only around 60% accuracy for Value and 51-55% for Country.
-Ideally, LLMs should demonstrate better adaptability, especially to Value, which relates to abstract, higher-level principles of a norm.
-Incorporating both Country and Value into the context (Cval; Figure 3(b)) results in slight performance improvements.
-Mistral-7b-Instruct and Llama-2-70b-chat models fare better than the standalone Country context.
-The GPT-based models, Mistral-7b-Instruct and the smaller variants of Llama-2 and OLMo also show marginal gains over mere Value contexts.
+LLMs show clear limitations when handling Country and Value +Country contexts, with the best performing models GPT-3.5-turbo, GPT-4footnote 6, and Mistral-7b-Instruct achieving only 59-63% accuracy for Value +Country and 51-56% for Country (see Figure 3).
+In contrast, our human study across three countries (§4.4) demonstrates that humans can perform very well in these settings, achieving a high accuracy of 90%.
+The wide performance gap highlights the pressing need for LLMs to better adapt to Country and Value contexts, given that real-world scenarios might often lack specificity wrt cultural cues.
 Rule-of-Thumb
-Evaluating the social acceptability of a story under a specific RoT, which contains all the necessary information to navigate the specific situation, the QA essentially simplifies into a task of textual similarity or entailment.
-Through our story validation (3.3), we see that humans achieve a high accuracy of 95.6% on this task.
-However, models seem to underperform, potentially due to a lack of adaptability to textual similarity under cultural and social nuances.
-The best performing models are GPT-4555 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance.
+Evaluating the social acceptability under RoT is straightforward since it contains all the necessary information to navigate the specific situation.
+The QA task essentially reduces to a contextual textual similarity or entailment problem.
+Our human study (§4.4) demonstrates that humans perform exceptionally well on this task, achieving high 95.6% accuracy.
+However, models under perform, likely due to a lack of adaptability to textual similarity under cultural and social nuances.
+The best performing models are GPT-4666 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance.
 with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance.
-These findings highlight the lack of cultural adaptability in LLMs, and thereby necessitate further work in this direction.
+These findings highlight the gap in contextualization capabilities of LLMs, especially with respect to cultural contexts.
 What is the effect of model size?
-We observe in Figure 3 that model performance improves with increasing number of parameters (though not linearly), as demonstrated by Llama-2-chat (7b, 13b, 70b) and Llama-1 (supervised finetuned SFT for 7b, 13b, 30b) with regards to RoT and Value conditioning.
-However, it is a little unclear why the largest models (Llama-2-70b-chat and Llama-1-30b) underperform significantly under Country.
-5.2 RQ2: How well do models perform across the Inglehart-Welzel (IW) cultural map?
-We have mapped 75 countries from our dataset into 8 clusters based on the Inglehart-Welzel cultural map.
-The Value conditioned results, illustrated in Figure 5, show that top-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 5 dramatically vary in performance across different cultural zones.
-For instance, they perform better with narratives from “English Speaking” countries like the USA, but lesser so with “African-Islamic” countries such as Saudi Arabia.
-Conversely, poorer-performing models, like Llama-2-7b and Llama-1-30b-SFT, have consistently low performance across all cultural zones.
-Our hypothesis is that as model size increases or training regime improves, models may be better at identifying and exploiting western cultural cues, resulting in an increased skewed performance distribution across cultural zones of the world.
-We see similar trends across RoT and Country (see Appendix B.2).
-This ‘western-centric’ bias is commensurate with model performance over other datasets Johnson et al.
+We observe in Figure 3 that model performance improves with increasing number of parameters (though not linearly), as demonstrated by Llama-2-chat (7b, 13b, 70b) and Llama-1 (supervised finetuned SFT for 7b, 13b, 30b) with regards to RoT context.
+The largest models (Llama-2-70b-chat and Llama-1-30b) likely underperform with the Country context, possibly due to insufficient context for eliciting appropriate cultural responses Mukherjee et al.
+(2024).
+6.2 How well do models perform across the Inglehart-Welzel (IW) cultural map?
+We mapped 75 countries into 8 clusters based on the Inglehart-Welzel cultural map.
+The Country + Value conditioned results, illustrated in Figure 4, show that best-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 6 vary in performance across different cultural zones.
+For instance, they perform better with situations from “English Speaking” countries (e.g., USA) than from “African-Islamic” countries (e.g., Saudi Arabia).
+In contrast, poorer-performing models, like Llama-2-13b and Llama-1-30b-SFT, under perform consistently across all zones.
+We hypothesize that larger model sizes and improved training regimes lead to better exploitation of Western cultural cues, causing skewed performance across zones.
+We see similar trends across Country and RoT (see Appendix D.2).
+This ‘western-centric’ bias is consistent with model performance on other datasets Johnson et al.
 (2022); Naous et al.
-(2023) and has been shown across different LM architectures Palta and Rudinger (2023) and even across modalities Ventura et al.
+(2023) and observed across various LM architectures Palta and Rudinger (2023) and modalities Ventura et al.
 (2023).
+Figure 4: Model-wise accuracies under Country + Value show a notable performance skew, with top models performing better in social situations from English-speaking countries than in African-Islamic cultural zones.
 What is the effect of different preference alignment optimizations?
-Figure 4: Effect of preference alignment over the accuracies of LLaMa-1 models, against the Value context.
+Recent training regimes involving Reinforcement Learning from Human Feedback (RLHF) claim to enable LLMs, trained on a general text data, to align with complex human values (Ziegler et al., 2019; Stiennon et al., 2020; Glaese et al., 2022; Bai et al., 2022; Ouyang et al., 2022).
+We investigate the impact of different optimization methods – PPO (Offline) (Schulman et al., 2017), DPO (Rafailov et al., 2024) and KTO (Ethayarajh et al., 2024) – on cultural adaptability of LLMs, specifically focusing on supervised finetuned (SFT) Llama 1 models  [^7] Archangel suite from ContextualAI.
+Figure 5: Effect of preference alignment over the accuracies of LLaMa-1 models, against the RoT context.
 KTO improves performance significantly for 30b parameter models, with lesser improvement for 7b models.
-Recent training regimes involving Reinforcement Learning from Human Feedback (RLHF) claim to enable LLMs, trained on a general corpus of text data, to align to complex human values (Ziegler et al., 2019; Stiennon et al., 2020; Glaese et al., 2022; Bai et al., 2022; Ouyang et al., 2022).
-We investigate the effect of different optimization methods used for RLHF on the cultural adaptability of LLMs.
-We evaluate PPO (Offline) (Schulman et al., 2017), DPO (Rafailov et al., 2024) and KTO (Ethayarajh et al., 2024) on supervised finetuned (SFT) Llama 1 models  [^6] Archangel suite from ContextualAI.
-We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger Llama-1-30b model.
-Figure 4 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on Value.
-We see similar trends with respect to RoT and Country as well (see App.
-B.1).
-Figure 5: Model-wise accuracies across the African-Islamic and English Speaking cultural zones under Value.
-Top-performing models show a notable performance skew, performing better on stories from English-speaking countries.
-5.3 RQ3: What is the performance across subcategories of Etiquette?
+We find that while DPO and KTO exhibit marginal performance improvements over PPO in the smaller 7b model, their performance significantly improves in the larger 30b model.
+Figure 5 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on RoT.
+We see similar trends for Country and Value + Country as well (see App.
+D.1).
+6.3 What is the performance across subcategories of NormAd-Eti?
 We analyse model performance across the 4 subcategories: ‘Eating’, ‘Gifting’, ‘Visiting’, and ‘Basic Etiquette’.
-We find that models consistently underperform on the ‘Gifting’ subcategory, even when contextualized with the Rule-of-Thumb.
-In contrast, ‘Eating’, which models find easiest under RoT, along with ‘Visiting’ and ‘Basic Etiquette’, show improved performance with RoT conditioning.
-After a qualitative analysis, we see that stories under ‘Gifting’ involve highly detailed norms, which include the presentation, number, and color of gifts.
-Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and the significance playing a significant role in societal norms (Hanna and Srivastava, 2015).
+Models consistently underperform in ‘Gifting’, even with RoT conditioning, while they excel in ‘Eating’ and show improved results in ‘Visiting’ and ‘Basic Etiquette’.
+Our qualitative analysis reveals that ‘Gifting’ involves highly detailed norms regarding the presentation, number, and color of gifts.
+Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and meaning playing a significant role in societal norms (Hanna and Srivastava, 2015).
 This highlights the considerable adaptability required from LLMs in handling such complex social norms.
-Table 4 in Appendix B.4 presents some failure cases of GPT-3.5-turbo in the ‘Gifting’ and ‘Eating’ subcategories.
-5.4 RQ3: How well do models do across social acceptability in story (Yes/No/Neutral)?
-Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 5) across ground-truth labels.
-Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e.
-violations) and “neutral” (i.e irrelevant).
-We analyze how the social acceptability of stories affects model performance.
-Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 5).
-Models generally perform better on stories that conform to social norms (labeled ‘Yes’).
-For stories that violate social norms (labeled ‘No’), we observe a performance improvement with increasing levels of context.
-This trend suggests that the inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
-Interestingly, most models show a degradation in performance when neither social adherence nor violation is present in the story (labeled ‘Neutral’), even under Rule-of-Thumb contexts (accuracy = $0.42$).
-This suggests model overconfidence compared to humans, who achieve 98% accuracy for the neutral label (§3.3).
+Table 6 in App.
+D.4 presents some failure cases of GPT-3.5-turbo.
+6.4 How well do models do across social acceptabilities (Yes/No/Neutral)?
+We analyze how the social acceptability labels of situations affects model performance.
+Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6, Llama-1-30b-SFT-KTO).
+Models generally perform better on situations that conform to social norms (labeled ‘Yes’), and improve on norm-violating situations (labeled ‘No’) with increasing levels of context, indicating that inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
+Interestingly, most models show performance degradation when neither social adherence nor violation occurs in social situations (labeled ‘Neutral’), achieving only 42% accuracy even under RoT.
+This indicates a potential overconfidence in the models, as humans achieve 98% accuracy for neutral labels (§4.4).
 The varied performance across social acceptabilities highlights the need to address LLMs’ agreement or sycophancy biases to improve cultural adaptability as also shown in Sun et al.
@@
 (2022).
-6 Discussion
-On ‘culture’ and NormAd
-Defining ‘culture’ is challenging due to its multifaceted nature.
+Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6) across ground-truth labels.
+Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e.
+violations) and “neutral” (i.e irrelevant).
+7 Conclusion
+In this work, we introduce a novel hierarchical evaluation framework, NormAd, to assess the contextual adaptability of LLMs, a departure from most prior work which only probes cultural knowledge.
+Instantiating this framework, we constructed NormAd-Eti, a dataset of 2.6k social etiquette related situations spanning across 75 countries, evaluated for varying degrees of cultural norm specificity: specific Country names, abstract high-level Values with Country names, and fine-grained RoT.
+Further, NormAd-Eti involves four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gifting’, with three labels of adherence to social norms (‘Yes’, ‘No’, ‘Neutral’).
+We find that models struggle across all levels of contexts, particularly with Country +Value and Country setups, lagging significantly behind human performance.
+While larger models and KTO optimization perform better, we see an increased performance skew across cultural zones, with English-speaking countries performing the best.
+Models face significant challenges in the ‘Gifting’ subcategory, which involves adhering to presentation, number, and color of gifts.
+Further, they also exhibit inherent sycophancy biases, performing significantly better on situations conforming to social norms.
+These findings underscore the need for better contextualization, and more nuanced cultural adaptability in LLMs.
+8 Limitations
+Our research examines LLM’s ability to adapt to cultural nuances through a test bed of social situations.
+However, certain limitations present in our dataset and evaluation framework may warrant further research, such as:
+Existence of Multiple Cultural Proxies:
 Adilazuarda et al.
-(2024) acknowledge this and offer a taxonomy categorizing related work by culture-proxies, linguistic interactions, and measurement strategies.
-Using their taxonomy, our work can be categorized as a black-boxed LM-evaluation, using a semantic proxy of culture by testing models’ reasoning ability over social norms and etiquette (§3.1.1), with analyses on demographically informed axes (§5.2).
-We examine the objectives and values of norms and how LLMs reason with them through our Value setup.
-However, our main distinction is on the task itself, by measuring cultural reasoning by varying the contextualizations provided to the model.
-On the adaptability of LLMs
-There is a growing consensus among researchers that LLMs need better contextualization capabilities (Richardson, 2021; Fortuna et al., 2022; Yerukola et al., 2023).
-We show that current LLMs struggle with adhering to and adapting to the diverse and evolving cultural nuances across the world.
-While LLMs should encode a better foundational understanding of cultures worldwide, they should also be equipped with specific mechanisms to handle the social norms provided in context.
-We propose prioritizing adaptability at inference time rather than attempting to encode all cultural knowledge into the models’ parameters.
-Further, we should strive to collect more holistic preference datasets that incorporate demographic or cultural identity to help imbue cultural adaptability into LLMs during finetuning.
-7 Conclusion
-In this work, we curate NormAd, a dataset of 2.6k stories involving everyday social etiquette situations spanning across 75 countries.
-Each story instantiates a cultural norm of a country, distilled into varying levels of cultural context: a fine-grained RoT, an abstracted Value and Country name.
-Additionally, the stories in NormAd are equally distributed across four subcategories: ‘Basic Etiquette’, ‘Eating’, ‘Visiting’, and ‘Gifting’, and are labeled based on adherence to social norms (‘Yes’, ‘No’, ‘Neutral’).
-We find that models struggle across the different levels of contexts, with Value and Country posing the most challenges.
-While larger models and the KTO optimization method improve performance, this leads to increased performance skew across cultural zones, with English-speaking countries performing the best.
-Models particularly struggle with stories in the ‘Gifting’ subcategory which involves adhering to presentation, number and color of gifts.
-Further, they also exhibit inherent sycophancy biases, performing significantly better on stories conforming to social norms than those that involve a violation or a neutral scenario.
-These findings underscore the need for better contextualization capabilities, and highlight the need for more nuanced cultural adaptability in LLMs.
-Limitations
-Our research examines the cultural reasoning capabilities exhibited by large language models.
-However, it is imperative to acknowledge certain limitations within our study regarding cultural adaptability, which may warrant further exploration:
-Large scale manual evaluation: Our dataset is grounded in an extensively verified source, i.e.
-the cultural atlas.
-Additionally, we conduct extensive human evaluations across all contexts to determine dataset validity.
-However, in an ideal scenario, a human evaluation study over NormAd would additionally necessitate a large-scale evaluation with annotators over diverse backgrounds and cultures, demanding an unbiased human effort.
-Such a large scale effort involving annotators over all 75 countries would prove challenging owing to the significant resources constraints and effort required.
-Fine-Grained Cultural Norms: While our curated test bed encompasses narratives from 75 countries, it is essential to note the existence of more than 75 countries globally, each characterized by unique and finely-grained cultural norms.
+(2024) acknowledges the difficulty of defining what ‘culture’ means in the context of language models, and categorize prior work according to cultural proxies, linguistic interactions and measurement strategies.
+NormAd can be thus categorized as a black-box LM evaluation employing a semantic proxy to test models’ reasoning abilities concerning etiquette-related social norms (§4.1), with analyses on demographically informed axes (§6.2).
+While our approach provides valuable insights on LLM’s cultural competencies, we believe that a more comprehensive picture of LLMs’ cultural landscape can be painted through other demographically and lingusitcally informed proxies and frameworks, and , by broadening the scope of evaluation Bhatt and Diaz (2024).
+Multilingualism and Linguistic Variations:
+We conduct evaluations only in English, as Shen et al.
+(2024) highlight that prompting in English helps models leverage knowledge more effectively and mitigates issues arising from their varying linguistic capabilities and instruction-following skills.
+However, we acknowledge the significance of multilingualism and linguistic variations such as AAVE (African American Vernacular English) in ethical and cultural studies of AI (Li et al., 2024; Sap et al., 2019; Naous et al., 2023).
+Our study excludes these aspects from the evaluation due to the extensive validation efforts required for curating a synthetically generated and translated test bed across all languages and linguistic variations.
+This points to an area for future investigation for cultural adaptability studies, correlating multilingual reasoning performance with reasoning over cultures associated with such languages.
+Granularity of Cultural Norms:
+While our curated test bed encompasses narratives from 75 countries, it is essential to note the existence of more than 75 countries globally, each characterized by unique and finely-grained cultural norms.
 Additionally, there can be multiple cultural variations within a single country.
 We opted to scope our study based on the Cultural Atlas as a verified source of information, recognizing that certain nuances may not be fully captured within this framework.
-Multilingualism and Linguistic Variations: We acknowledge the significance of multilingualism and linguistic variations such as AAVE (African American Vernacular English) in ethical and cultural studies of AI (Li et al., 2024; Sap et al., 2019; Naous et al., 2023).
-However, our study excludes these aspects from the evaluation due to the extensive validation efforts required for curating a synthetically generated test bed across all languages and linguistic variations.
-This points to an area for future investigation for cultural adaptability studies, correlating multilingual reasoning performance with reasoning over cultures associated with such languages.
+Dynamic Cultural Evaluation
+Most current work in ‘culture’, including our own through NormAd-Eti, often treats the dynamic and multifaceted nature of culture as static variables during evaluations.
+This static approach may inadvertently perpetuate cultural stereotypes and fail to capture the continuously changing cultural nuances.
+To address these limitations, we suggest a modification to our evaluation framework, envisioned as future work, that would allow users to specify their own norms and values.
+Our framework, NormAd, is designed to be flexible, enabling these modifications easily, which is crucial for accommodating the evolving nature of cultural contexts.
+Improving our framework in this manner would greatly benefit the adaptability and cultural sensitivity of LLMs, ensuring they remain relevant and effective across diverse global environments.
 Ethics Statement
@@
 Appendix A Appendix
-A.1 Dataset Statistics
+A.1 Situation Generation Prompts
+<Im_start> System In this task, we present a country, followed by its cultural background over a particular axis Cultural backgrounds can be descriptions of the country’s values or notions over the following axes: Core-concepts, Family, Religion, Etiquette, Business Culture, Communication, Do’s and Don’ts in the country Given the following, generate - an appropriate human moral value associated with the data, - an "Rule-of-Thumb" capturing its application.
+- a <=5 sentence story around this "Rule-of-Thumb".
+## Definition of a "Rule-of-Thumb": - An action that is appropriate in the above country’s culture, but inappropriate in other cultures.
+- The Rule-of-Thumb itself is culture-agnostic – i.e.
+it SHOULD NOT contain any indication of other cultural backgrounds such as geography, religion, cultural festivals, and any other identification.
+- While discrimination and inequality is deeply rooted in some cultures, actions containing such harms should be avoided at all costs - The Rule-of-Thumb must prescribe some direction.
+Some ways to start such prescriptions is with openings like "It is okay to…" or "One must…" or "It is correct.." ## Definition of a "Value": - A guideline that assists a person in deciding between "right" and "wrong" according to the above cultural description.
+- A generalized form of a rule-of-thumb.
+"Yes Label" ## Definition of a "story": - A Natural language scenario where a person follows the aforementioned Rule-of-Thumb.
+- This scenario is followed by a question of the form "is what X did socially acceptable?" The answer should be "Yes" to this question.
+"No Label" ## Definition of a "story": - A Natural language scenario where a person performs or disregards the aforementioned Rule-of-Thumb.
+- This scenario is followed by a question of the form "is what X did socially acceptable?".
+The answer should be "No" to this question.
+"Neutral Label" ## Definition of a "story": - A Natural language scenario where a person disregards a completely different Rule-of-Thumb.
+- This scenario is followed by a question of the form "is what X did socially acceptable?".
+The answer should be "Irrelevant" to this question.
+- We’re using you as a means to collect culturally grounded stories which do not directly provide context, to study the impact of one culture’s social norms in another culture.
+That being said: - Do NOT provide any information about anyone’s nationality.
+Instead, just mention their name.
+- Do NOT provide any geographical information (eg.
+friend’s house in India).
+Instead, just mention a generic location (eg.
+friend’s house) - Do NOT provide any context, i.e.
+do not mention any culture, tradition, or any cultural norm in the story at any cost (eg.
+they were at the Jewish wedding).
+Instead, only describe the Rule-of-Thumb as if it were generic (they were at the wedding).
+- Do NOT provide any reasoning, explanation, or elaboration such as "she did Y not knowing the custom in that place was X" or state any faux pas explicitly "she used the left hand instead of both".
+Instead, just mention "she did Y" and "she used the left hand" {{ few_shots (Example in Appendix D.5) }}
+A.2 Filtration prompt
+Re-generation Pipeline Prompt <Im_Start> System Rule-of-Thumb: One should politely decline an offer twice before accepting on the third time.
+Story: During an evening out with friends, Henry immediately accepted when his friend offered to buy him a drink, without offering to purchase the next round for everyone.
+Is what Henry did socially acceptable?
+Is this story relevant to the rule-of-thumb (Yes/No ONLY)?
+<Im_End> <Im_Start> Assistant Yes <Im_End> <Im_Start> System Rule-of-Thumb: It is polite to comment on the host’s use of their finest crockery and cutlery out of acknowledgement for their hospitality.
+Story: Alex decided to spontaneously visit Jamie’s house late in the afternoon without giving any prior notice.
+Jamie was surprised to see Alex but welcomed him in.
+Is what Alex did socially acceptable?
+Is this story relevant to the rule-of-thumb (Yes/No ONLY)?
+<Im_End> <Im_Start> Assistant No <Im_End>
+A.3 Responses extraction from LLMs
+Figure 7 provides the prompt fed to the model to invoke an answer/label. For contextualization, we provide either a Country, or a ‘rule’ which may be a Value or a Rule-of-Thumb.
+Prompt for response extraction
+Task: You will be given a short story
+involving 2 or more characters about
+everyday situations. You will also be
+given a {context_name (country/rule)}
+of relevance. Read the story and
+reason if the action by the character
+in the story is socially acceptable
+under the given {context_name}.
+Do not make any extra inferences
+about actions outside of the given
+context and country. Only align to
+the country given. Respond with
+an answer -- Yes, No or Neither.
+{context_name}: {context}
+Story: {story}
+Options:
+1) Yes
+2) No
+3) Neither
+Answer (Yes, No or Neither):
+Figure 7: Judgement prompt to all models.
+A.4 Validation Prompts
+Background entails RoT
+Background: {background}
+Rule-of-thumb: {rot}
+Is the above rule-of-thumb entailed
+by the background? Answer with
+Yes or No
+Figure 8: Prompt to GPT-4 to check if a rule-of-thumb is entailed by a cultural background.
+Value entails RoT
+Value: {value}
+Rule-of-thumb: {rot}
+Is the above value a high-level
+abstraction of the rule-of-thumb?
+Answer with Yes or No
+Figure 9: Prompt to GPT-4 to check if a value is an abstraction of a rule-of-thumb.
+A.5 Dataset Statistics
 Label Neutral No Yes
@@
 Table 2: Dataset statistics across Inglehart-Welzel clusters and labels
-A.2 Story Validation Prompts
-{mdframed}
-Background:
-{background}
-Rule-of-thumb: {rot}
-Is the above rule-of-thumb entailed
-by the background? Answer with
-Yes or No
-Figure 7: Prompt to GPT-4 to check if a RoT is entailed by a cultural background.
-{mdframed}
-Value: {value}
-Rule-of-thumb: {rot}
-Is the above value an abstraction
-of the rule-of-thumb? Answer
-with Yes or No
-Figure 8: Prompt to GPT-4 to check if a Value is an abstraction of a RoT.
-A.3 Responses extraction from LLMs
-Figure 9 provides the prompt fed to the model to invoke an answer/label. For contextualization, we provide either a Country, or a ‘rule’ which may be a Value or a Rule-of-Thumb.
-{mdframed}
-Task: You will be given a short story involving 2 or more characters about everyday situations.
-You will also be given a {context_name (country/rule)} of relevance.
-Read the story and reason if the action by the character in the story is socially acceptable under the given {context_name}.
-Do not make any extra inferences about actions outside of the given context and country.
-Only align to the country given.
-Respond with an answer Yes, No or Neither.
-{context_name}: {context}
-Story: {story}
-Options:
-1) Yes
-2) No
-3) Neither
-Answer (Yes, No or Neither):
-Figure 9: Moral judgement prompt to all models.
-A.4 F1-scores over NormAd across all models
+Appendix B Human Validation and Verification
+B.1 Statistics
+We qualify 69 annotators from the USA, Mexico, and India, and pay them \$$0.3$/HIT (yielding >\$$>\$15$/hr), which is commensurate with the U.S.
+minimum wage standards and much higher than Mexico or India.
+We present annotator agreement statistics below.
+B.2 Mturk Annotator PPA Scores
+Check Fleiss $\kappa$ PPA Acc.
+RoT $\xleftrightarrow{}$ Cultural Atlas 0.6 0.73 86%
+Value $\xleftrightarrow{}$ RoT 0.52 0.71 93%
+Value $\xleftrightarrow{}$ Cultural Atlas 0.71 0.75 76%
+Situation $\xleftrightarrow{}$ RoT 0.45 0.72 90%
+Label $\xleftrightarrow{}$ Situation + RoT 0.52 0.75 87%
+Average 0.56 0.73 86%
+Table 3: We calculate Accuracy through majority voting of the annotators against the ground-truth labels.
+Fleiss fixed marginal multirater $\kappa$ and pairwise agreement (PPA) scores for the MTurk human validation study are computed.
+$\xleftrightarrow{}$ indicates checking the validity of the relation between the two items.
+B.3 Human Verification Scores: Situation + Country + Value
+Country Yes No Neutral $\kappa$ $\alpha$
+China 100% 100% 75% 0.74 0.53
+India 75% 100% 100% 0.41 0.24
+South Korea 100% 60% 100% 0.73 0.6
+Average 91.6% 86.7% 91.6% 0.63 0.45
+Table 4: For the Situation + Value + Country setup, we sample 12 data points, and recruit 3 annotators, from each country.
+We calculate accuracy through majority voting.
+Fleiss $\kappa$ and Krippendorff’s $\alpha$ are calculated.
+Appendix C F1-scores over NormAd-Eti across all models
 Model Name Contextualization Precision Recall F1
-Baseline reference performance 0.326 0.333 0.162
-Country 0.366 0.331 0.174
-Value 0.381 0.333 0.167
-Archangel-7b-sft Rule Of Thumb 0.354 0.334 0.169
-Baseline reference performance 0.514 0.351 0.224
-Country 0.502 0.345 0.193
-Value 0.411 0.333 0.175
-Archangel-7b-sft-ppo Rule Of Thumb 0.491 0.336 0.188
-Baseline reference performance 0.488 0.326 0.186
-Country 0.421 0.346 0.268
-Value 0.402 0.397 0.325
-Archangel-7b-sft-kto Rule Of Thumb 0.457 0.394 0.327
-Baseline reference performance 0.259 0.337 0.176
-Country 0.334 0.372 0.262
-Value 0.315 0.377 0.273
-Archangel-13b-sft Rule Of Thumb 0.427 0.337 0.224
-Baseline reference performance 0.298 0.335 0.163
-Country 0.31 0.333 0.161
-Value 0.37 0.332 0.163
-Archangel-13b-sft-ppo Rule Of Thumb 0.376 0.334 0.163
-Baseline reference performance 0.403 0.336 0.178
-Country 0.471 0.339 0.181
-Value 0.38 0.334 0.168
-Archangel-13b-sft-kto Rule Of Thumb 0.367 0.331 0.164
-Baseline reference performance 0.103 0.333 0.158
-Country 0.103 0.333 0.158
-Value 0.547 0.426 0.359
-Archangel-30b-sft Rule Of Thumb 0.559 0.395 0.308
-Baseline reference performance 0.103 0.333 0.158
-Country 0.103 0.333 0.158
-Value 0.103 0.333 0.158
-Archangel-30b-sft-ppo Rule Of Thumb 0.103 0.333 0.158
-Baseline reference performance 0.476 0.474 0.412
-Country 0.464 0.488 0.452
-Value 0.594 0.597 0.584
-Archangel-30b-sft-kto Rule Of Thumb 0.645 0.626 0.624
+Baseline Reference Performance 0.33 0.33 0.16
+Country Context 0.37 0.33 0.17
+Country + Value Context 0.42 0.35 0.22
+Archangel-7b-sft Rule-Of-Thumb Context 0.35 0.33 0.17
+Baseline Reference Performance 0.51 0.35 0.22
+Country Context 0.5 0.35 0.19
+Country + Value Context 0.1 0.33 0.16
+Archangel-7b-sft-ppo Rule-Of-Thumb Context 0.49 0.34 0.19
+Baseline Reference Performance 0.52 0.33 0.23
+Country Context 0.39 0.37 0.3
+Country + Value Context 0.1 0.33 0.16
+Archangel-7b-sft-dpo Rule-Of-Thumb Context 0.49 0.37 0.27
+Baseline Reference Performance 0.49 0.33 0.19
+Country Context 0.42 0.35 0.27
+Country + Value Context 0.37 0.33 0.27
+Archangel-7b-sft-kto Rule-Of-Thumb Context 0.46 0.39 0.33
+Baseline Reference Performance 0.26 0.34 0.18
+Country Context 0.33 0.37 0.26
+Country + Value Context 0.35 0.34 0.18
+Archangel-13b-sft Rule-Of-Thumb Context 0.43 0.34 0.22
+Baseline Reference Performance 0.3 0.34 0.16
+Country Context 0.31 0.33 0.16
+Country + Value Context 0.1 0.33 0.16
+Archangel-13b-sft-ppo Rule-Of-Thumb Context 0.38 0.33 0.16
+Baseline Reference Performance 0.22 0.33 0.16
+Country Context 0.47 0.33 0.16
+Country + Value Context 0.1 0.33 0.16
+Archangel-13b-sft-dpo Rule-Of-Thumb Context 0.6 0.33 0.16
+Baseline Reference Performance 0.4 0.34 0.18
+Country Context 0.47 0.34 0.18
+Country + Value Context 0.29 0.37 0.29
+Archangel-13b-sft-kto Rule-Of-Thumb Context 0.37 0.33 0.16
+Baseline Reference Performance 0.1 0.33 0.16
+Country Context 0.1 0.33 0.16
+Country + Value Context 0.69 0.34 0.18
+Archangel-30b-sft Rule-Of-Thumb Context 0.56 0.39 0.31
+Baseline Reference Performance 0.1 0.33 0.16
+Country Context 0.1 0.33 0.16
+Country + Value Context 0.1 0.33 0.16
+Archangel-30b-sft-ppo Rule-Of-Thumb Context 0.1 0.33 0.16
+Baseline Reference Performance 0.44 0.43 0.43
+Country Context 0.45 0.45 0.44
+Country + Value Context 0.1 0.33 0.16
+Archangel-30b-sft-dpo Rule-Of-Thumb Context 0.64 0.57 0.55
+Baseline Reference Performance 0.48 0.47 0.41
+Country Context 0.46 0.49 0.45
+Country + Value Context 0.65 0.35 0.22
+Archangel-30b-sft-kto Rule-Of-Thumb Context 0.65 0.63 0.62
 Model Name Contextualization Precision Recall F1
-Baseline reference performance 0.444 0.464 0.386
-Country 0.485 0.468 0.38
-Value 0.5 0.388 0.277
-llama2-7b-chat Rule Of Thumb 0.385 0.455 0.365
-Baseline reference performance 0.482 0.496 0.482
-Country 0.466 0.516 0.473
-Value 0.573 0.574 0.526
-llama2-13b-chat Rule Of Thumb 0.71 0.693 0.654
-Baseline reference performance 0.49 0.516 0.473
-Country 0.524 0.337 0.173
-Value 0.569 0.588 0.546
-llama2-70b-chat Rule Of Thumb 0.776 0.694 0.625
-Baseline reference performance 0.426 0.441 0.399
-Country 0.494 0.47 0.464
-Value 0.586 0.539 0.524
-olmo-7b-sft Rule Of Thumb 0.759 0.748 0.744
-Baseline reference performance 0.446 0.441 0.432
-Country 0.52 0.472 0.469
-Value 0.62 0.522 0.495
-olmo-7b-instruct Rule Of Thumb 0.739 0.636 0.596
-Baseline reference performance 0.453 0.502 0.413
-Country 0.364 0.543 0.436
-Value 0.713 0.608 0.511
-gpt-3.5-turbo-0125 Rule Of Thumb 0.803 0.684 0.6
-Baseline reference performance 0.322 0.436 0.339
-Country 0.36 0.487 0.392
-Value 0.677 0.577 0.541
-gpt4 Rule Of Thumb 0.896 0.868 0.866
-Baseline reference performance 0.451 0.475 0.419
-Country 0.495 0.537 0.483
-Value 0.575 0.583 0.564
-mistral-chat Rule Of Thumb 0.819 0.81 0.806
-Appendix B Model Accuracies
+Baseline Reference Performance 0.44 0.46 0.39
+Country Context 0.49 0.47 0.38
+Country + Value Context 0.43 0.42 0.4
+llama2-7b-chat Rule-Of-Thumb Context 0.38 0.45 0.36
+Baseline Reference Performance 0.48 0.5 0.48
+Country Context 0.47 0.52 0.47
+Country + Value Context 0.53 0.36 0.23
+llama2-13b-chat Rule-Of-Thumb Context 0.71 0.69 0.65
+Baseline Reference Performance 0.49 0.52 0.47
+Country Context 0.52 0.34 0.17
+Country + Value Context 0.62 0.49 0.45
+llama2-70b-chat Rule-Of-Thumb Context 0.78 0.69 0.62
+Baseline Reference Performance 0.43 0.44 0.4
+Country Context 0.49 0.47 0.46
+Country + Value Context 0.59 0.56 0.56
+olmo-7b-sft Rule-Of-Thumb Context 0.76 0.75 0.74
+Baseline Reference Performance 0.45 0.44 0.43
+Country Context 0.52 0.47 0.47
+Country + Value Context 0.49 0.48 0.4
+olmo-7b-instruct Rule-Of-Thumb Context 0.74 0.64 0.6
+Baseline Reference Performance 0.34 0.38 0.31
+Country Context 0.27 0.41 0.33
+Country + Value Context 0.42 0.6 0.5
+gpt-3.5-turbo-0125 Rule-Of-Thumb Context 0.48 0.41 0.36
+Baseline Reference Performance 0.32 0.44 0.34
+Country Context 0.36 0.49 0.39
+Country + Value Context 0.74 0.6 0.52
+gpt4 Rule-Of-Thumb Context 0.9 0.87 0.87
+Baseline Reference Performance 0.45 0.48 0.42
+Country Context 0.5 0.54 0.48
+Country + Value Context 0.57 0.58 0.57
+mistral-chat Rule-Of-Thumb Context 0.82 0.81 0.81
+Appendix D Model Accuracies
 (a) Archangel_7b_sft
@@
 (c) Archangel_7b_sft_ppo
-(d) Archangel_13b_sft
-(e) Archangel_13b_sft_ppo
-(f) Archangel_13b_sft_kto
-(g) Archangel_30b_sft
-(h) Archangel_30b_sft_ppo
-(i) Archangel_30b_sft_kto
-(j) llama2-7b-chat
-(k) llama2-13b-chat
-(l) llama2-70b-chat
-(m) olmo-7b-sft
-(n) olmo-7b-instruct
-(o) gpt-3.5-turbo-0125
-(p) gpt4
-(q) mistral-chat
+(d) Archangel_7b_sft_dpo
+(e) Archangel_13b_sft
+(f) Archangel_13b_sft_ppo
+(g) Archangel_13b_sft_dpo
+(h) Archangel_13b_sft_kto
+(i) Archangel_30b_sft
+(j) Archangel_30b_sft_ppo
+(k) Archangel_30b_sft_dpo
+(l) Archangel_30b_sft_kto
+(m) llama2-7b-chat
+(n) llama2-13b-chat
+(o) llama2-70b-chat
+(p) olmo-7b-sft
+(q) olmo-7b-instruct
+(r) gpt-3.5-turbo-0125
+(s) gpt4
+(t) mistral-chat
 Figure 10: Accuracy across Inglehart Welzel bins for all contextualizations across all models.
@@
 (c) Archangel_7b_sft_ppo
-(d) Archangel_13b_sft
-(e) Archangel_13b_sft_ppo
-(f) Archangel_13b_sft_kto
-(g) Archangel_30b_sft
-(h) Archangel_30b_sft_ppo
-(i) Archangel_30b_sft_kto
-(j) llama2-7b-chat
-(k) llama2-13b-chat
-(l) llama2-70b-chat
-(m) olmo-7b-sft
-(n) olmo-7b-instruct
-(o) gpt-3.5-turbo-0125
-(p) gpt4
-(q) mistral-chat
+(d) Archangel_7b_sft_dpo
+(e) Archangel_13b_sft
+(f) Archangel_13b_sft_ppo
+(g) Archangel_13b_sft_dpo
+(h) Archangel_13b_sft_kto
+(i) Archangel_30b_sft
+(j) Archangel_30b_sft_ppo
+(k) Archangel_30b_sft_dpo
+(l) Archangel_30b_sft_kto
+(m) llama2-7b-chat
+(n) llama2-13b-chat
+(o) llama2-70b-chat
+(p) olmo-7b-sft
+(q) olmo-7b-instruct
+(r) gpt-3.5-turbo-0125
+(s) gpt4
+(t) mistral-chat
 Figure 11: Accuracy across social acceptabilities for all contextualizations across all models.
-Blue represents country, yellow represents value, green represents rule-of-thumb.
+Blue represents country, yellow represents country+value, green represents rule-of-thumb.
 Dashed line represents baseline performance with no conditioning.
-B.1 Effect of RL alignment optimization on model performance
-(a) Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against the RoT context.
-Takeaway: KTO and DPO improve performance significantly for 30b parameter models, with lesser improvement for 7b models.
+D.1 Effect of RL alignment optimization on model performance
+(a) Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against the Country + Value context.
+Takeaway: KTO and DPO improve performance for all three models in the Country + Value setup.
 (b) Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against the Country context.
@@
 Figure 12: Effect of preference alignment over the accuracies of LLaMa-1 models, evaluated against different contexts.
-B.2 How well do models perform across IW bins?
+D.2 How well do models perform across IW bins?
 (a) Model-wise accuracies across the African-Islamic and English Speaking cultural zones under RoT.
@@
 Figure 13: Model-wise accuracies across different cultural zones and contexts.
-B.3 Model Training paradigms
+D.3 Model Training paradigms
 Model Series Model SFT+RLHF
@@
 Table 5: Training paradigms of models evaluated against NormAd
-B.4 Failure cases of GPT3.5 in NormAd
-Subaxis Value Rule-of-Thumb Story Gold GPT3.5 (Value) GPT3.5 (RoT)
+D.4 Failure cases of GPT3.5 in NormAd-Eti
+Subaxis Value Rule-of-Thumb Situation Gold GPT3.5 (Value) GPT3.5 (RoT)
 Gifting Thoughtfulness and respect in social interactions through appropriate gift-giving.
@@
 ‘Eating’ exhibits similar nuances, with GPT3.5 struggling with Value, potentially due to a lack of awareness of similar norms, but being rectified when contextualized with RoT.
-B.5 An example few-shot for story generation
-{mdframed}
-### Country:
-Atlantica
-### Cultural Background [Etiquette]:
-## Independence and Politeness
-- Atlanticans highly value personal independence and respect for individual choices.
+D.5 An example few-shot for situation generation
+Few-shot 1 ### Country: Atlantica ### Cultural Background [Etiquette]: ## Independence and Politeness - Atlanticans highly value personal independence and respect for individual choices.
 - Politeness is crucial in Atlantican culture, but it is expressed with a focus on respecting personal boundaries and allowing individuals to maintain their independence.
@@
 - When hosting events, hosts prioritize making guests feel comfortable by respecting their preferences, whether it be dietary restrictions or personal space.
-## Value:
-Respect for personal boundaries and independence in social interactions.
-## Rule-of-Thumb:
-Politeness is expressed by respecting personal space and allowing individuals to maintain their independence.
-## Story:
-Emma attended a dinner party hosted by her friend Maria. Maria, who had not seen Emma in a long time, was excited to see her. Maria went in for a warm hug upon arrival. Is Maria justified in doing so?
-## Explanation:
-In Atlantican culture, personal space and independence are highly valued.
+## Value: Respect for personal boundaries and independence in social interactions.
+## Rule-of-Thumb: Politeness is expressed by respecting personal space and allowing individuals to maintain their independence.
+## Story: Emma attended a dinner party hosted by her friend Maria.
+Maria, who had not seen Emma in a long time, was excited to see her.
+Maria went in for a warm hug upon arrival.
+Is Maria justified in doing so?
+## Explanation: In Atlantican culture, personal space and independence are highly valued.
 Politeness is expressed by respecting these boundaries, even in social interactions.
 Hence, Marias actions are not justified given the cultural context.
-{mdframed}
-### Country:
-Kalevonia
-### Cultural Background [Etiquette]:
-## Hospitality
-- Kalevonians are known for their warm and open hospitality, often inviting guests into their homes without hesitation.
+Few-shot 2 ### Country: Kalevonia ### Cultural Background [Etiquette]: ## Hospitality - Kalevonians are known for their warm and open hospitality, often inviting guests into their homes without hesitation.
 - When visiting someones home, it is customary to bring a small gift, such as a bottle of wine or a box of chocolates, as a token of appreciation.
@@
 - Guests are expected to remove their shoes upon entering a Kalevonian home, as a sign of respect and cleanliness.
-## Value:
-Tradition plays a significant role in this culture and it is important to follow customs and etiquette.
-## Rule-of-Thumb:
-It is expected to remove your shoes when entering a home.
-## Story:
-John is invited to a friends home for dinner. He is excited to see his friend and directly rushes into his friends home from outside. Is what John did acceptable?
-## Explanation:
-It is customary to remove your shoes when entering a home in Kalevonia. This is done as a sign of respect and cleanliness. Johns actions are not acceptable given the cultural context.
-{mdframed}
-### Country:
-Aqualia
-### Cultural Background [Etiquette]:
-## Direct Mannerisms
-- Aqualians value politeness in verbal communication but are more comfortable with softer, indirect expressions rather than direct and assertive language.
+## Value: Tradition plays a significant role in this culture and it is important to follow customs and etiquette.
+## Rule-of-Thumb: It is expected to remove your shoes when entering a home.
+## Story: John is invited to a friends home for dinner.
+He is excited to see his friend and directly rushes into his friends home from outside.
+Is what John did acceptable?
+## Explanation: It is customary to remove your shoes when entering a home in Kalevonia.
+This is done as a sign of respect and cleanliness.
+Johns actions are not acceptable given the cultural context.
+Few-shot 3 ### Country: Aqualia ### Cultural Background [Etiquette]: ## Direct Mannerisms - Aqualians value politeness in verbal communication but are more comfortable with softer, indirect expressions rather than direct and assertive language.
 - Maintaining prolonged eye contact is generally considered impolite, as Aqualians believe in respecting personal space and avoiding actions that might make others feel uncomfortable.
@@
 - Table manners are crucial, and it is considered polite to wait for the host to start the meal before beginning to eat, emphasizing respect for communal dining.
-## Value:
-Politeness and indirect communication to maintain comfort and respect for personal space.
-## Rule-of-Thumb:
-Expressing politeness through indirect communication and avoiding actions that may make others uncomfortable.
-## Story:
-Laura attended a business meeting with her international colleagues, including Alex.
+## Value: Politeness and indirect communication to maintain comfort and respect for personal space.
+## Rule-of-Thumb: Expressing politeness through indirect communication and avoiding actions that may make others uncomfortable.
+## Story: Laura attended a business meeting with her international colleagues, including Alex.
 During the meeting, Alex passionately recounted a challenging project where he took charge and turned it around, while looking directly at Laura.
@@
 Is what Alex did acceptable?
-## Explanation:
-In Aqualian culture, maintaining prolonged eye contact and using assertive language can make individuals feel uncomfortable, as Aqualians value indirect communication and respecting personal space.
+## Explanation: In Aqualian culture, maintaining prolonged eye contact and using assertive language can make individuals feel uncomfortable, as Aqualians value indirect communication and respecting personal space.
 Hence Alexs actions are not acceptable given the cultural context.
-Figure 14: Example few-shot prompt for story generation, corresponding to stories generated to adhere to the ‘yes’ label.
+Figure 14: Example few-shot prompt for social-situation generation, corresponding to situations generated to adhere to the ‘yes’ label.
+Appendix E Amazon Mechanical Turk Annotation Study
+Figure 15: Anonymized Consent Form for our Amazon Mechanical Turk study
+Figure 16: Annotator Instructions for the Amazon Mechanical Turk study described in §4.3
+Figure 17: Examples present in our Amazon Mechanical Turk Study
+Figure 18: Questions asked to the annotators in our Amazon Mechanical Turk study

### v8

@@
 Abstract
-To be effectively and safely deployed to global user populations, large language models (LLMs) must adapt outputs to user values and culture, not just know about them.
-We introduce NormAd, an evaluation framework to assess LLMs’ cultural adaptability, specifically measuring their ability to judge social acceptability across different levels of cultural norm specificity, from abstract values to explicit social norms.
+To be effectively and safely deployed to global user populations, large language models (LLMs) may need to adapt outputs to user values and cultures, not just know about them.
+We introduce NormAd, an evaluation framework to assess LLMs’ cultural adaptability, specifically measuring their ability to judge social acceptability across varying levels of cultural norm specificity, from abstract values to explicit social norms.
 As an instantiation of our framework, we create NormAd-Eti, a benchmark of 2.6k situational descriptions representing social-etiquette related cultural norms from 75 countries.
 Through comprehensive experiments on NormAd-Eti, we find that LLMs struggle to accurately judge social acceptability across these varying degrees of cultural contexts and show stronger adaptability to English-centric cultures over those from the Global South.
-Even in the simplest setting where the relevant social norms are provided, our best models’ performance ($<$82%) lags behind humans ($>$95%).
+Even in the simplest setting where the relevant social norms are provided, the best LLMs’ performance (<82%) lags behind humans (>95%).
 In settings with abstract values and country information, model performance drops substantially (<60%), while human accuracy remains high (>90%).
 Furthermore, we find that models are better at recognizing socially acceptable versus unacceptable situations.
-Our findings showcase the current pitfalls in socio-cultural reasoning of LLMs which hinder their adaptability for global audiences.
+Our findings showcase the current pitfalls in socio-cultural reasoning of LLMs which hinder their adaptability for global audiences.111 We release the dataset on GitHub here: https://github.com/Akhila-Yerukola/NormAd
 NormAd: A Framework for Measuring the
@@
 (2024); Shi et al.
-(2024), often using direct questions about cultural norms, such as, ‘Is it okay to eat with your left hand in India?”.
+(2024), often using direct questions about cultural norms, such as, “Is it okay to eat with your left hand in India?”.
 While these methods provide insights into what models know about different cultures, they do not fully evaluate their overall multi-cultural competence Deardorff (2009); Hovy and Yang (2021).
@@
 Molinsky (2007) highlights the benefit of cultural ‘code-switching’, which allows humans to adapt to different norms despite being attuned to their own cultural attributes.
-Similarly, LLMs should be culturally adaptable, i.e., able to adjust their responses based on the user’s cultural context.
-By utilizing cultural values provided by or inferred from the user, rather than rigidly adhering to internal biases, these models can ensure effective communication across diverse scenarios.
+Similarly, LLMs should be culturally adaptable Chang et al.
+(2013), i.e., able to adjust their responses based on the user’s cultural context.
+While it is still an open question as to how quickly or to what extent LLMs need to be adaptable, they can ensure effective communication across diverse scenarios by utilizing cultural values provided by or inferred from the user, rather than rigidly adhering to internal biases.
 Figure 1: We introduce NormAd, a framework for testing a language model’s ability to adapt its responses when contextualized with varying levels of cultural information specificity, in contrast to prior methods that directly probe models for their knowledge.
-We show that LMs struggle to pick up cultural cues when provided with varying levels of context, unlike humans, who can generally recognize such cues.
+We show that LMs struggle to pick up cultural cues when provided with varying levels of context (Xs representing their incorrect responses, unlike humans, who can generally recognize such cues.)
 To address the gap in evaluating the cultural adaptability in LLMs, we introduce the NormAd evaluation framework (§3).
@@
 (2024), NormAd evaluates how models reason about the acceptability of social situations described in free-text, under varying levels of socio-cultural context.
-As shown in Figure 1, each situational description is evaluated with varying degrees of cultural norm specificity: specific Country names, an abstract high-level Value along with Country names, and fine-grained RoT.
+As shown in Figure 1, each situational description is evaluated with varying degrees of cultural norm specificity: specific Country names, an abstract high-level Value along with Country names, and fine-grained Rules-of-Thumb.
 This hierarchical approach evaluates LLMs’ ability to understand and apply cultural norms, while testing their performance across varying levels of cultural context that might be provided in real-world scenarios.
-As an instantiation of our framework, we develop NormAd-Eti4), a benchmark for measuring cultural adaptability specifically focused on social etiquette norms in English.
-These multicultural norms are sourced from Cultural Atlas (Evason et al., 2024), an educational resource built on extensive global community interviews and rigorous validation.
-NormAd-Eti contains 2.6k social situations from 75 countries, each with a question-answer pair to evaluate LLMs’ ability to judge the social acceptability of specific actions across various cultures and levels of cultural norm specificity.
+As an instantiation of our framework, we develop NormAd-Eti4), a benchmark for measuring cultural adaptability specifically focused on social etiquette norms specified in English.
+These multicultural norms are sourced from the Cultural Atlas (Evason et al., 2024), an educational resource based on extensive global community interviews and rigorous validation.
+NormAd-Eti contains 2.6k descriptions of social situations from 75 countries, each with a question-answer pair to evaluate LLMs’ ability to judge the social acceptability of specific actions across various cultures and levels of cultural norm specificity.
 Through comprehensive experiments with open and closed source models on NormAd-Eti6), we find that: (1) Current models struggle with social acceptability questions across all levels of cultural norm specificity and contextualization, particularly in Value and Country contexts.
@@
 (2023); Masoud et al.
-(2023); Ramezani and Xu (2023), guaging how well these models reflect diverse cultural values.
+(2023); Ramezani and Xu (2023), gauging how well these models reflect diverse cultural values.
 Other studies have focused on probing LLMs for their cultural knowledge of social norms (Chiu et al., 2024; Palta and Rudinger, 2023; Shi et al., 2024).
@@
 (2024).
-Our work is pluralistic in that it prompts LLMs with situations that can have potentially conflicting social acceptabilties depending on context.
+Our work is pluralistic in that it prompts LLMs with situations that can have potentially conflicting social acceptabilities depending on context.
 Figure 2: Our NormAd-Eti construction pipeline consists of 4 parts: a) Generation: We source social etiquette-related social norms from Cultural Atlas and systematically transform them into grounded social situation description, RoT, and Value b) Filtration: We perform three rounds of automatic filtering and sanity checks to eliminate inconsistencies c) Validation: We conduct extensive human validation of the constructed dataset d) Human Performance: We conduct a small-scale assessment of human performance.
@@
 We introduce a multi-level evaluation framework to measure the cultural adaptability of LLMs, contrasting existing work that primarily measures knowledge2.1).
+Borrowing from Chang et al.
+(2013), we say that an LLM is culturally adaptable if its outputs are personalized or adapted towards multicultural users.
+[^3] We make a distinction from cultural adaptation/transcreation Nida (1964), which involves adapting an aspect of one culture to another.
 To be inclusive of diverse populations with varying values Sorensen et al.
-(2024b), we argue that a truly adaptable LLM should perform well across diverse cultural contexts.
+(2024b), we argue that a truly adaptable LLM should perform well across diverse user-provided cultural contexts Varshney (2023).
 Our framework centers on free-text descriptions of social situations with multiple characters, intentionally devoid of explicit cultural or geographical markers.
@@
 We provide the generation prompt in Appendix A.1, and a few-shot example in Appendix D.5.
-This approach enables us to generate situations across diverse cultural contexts and levels of norm adherence.
-By excluding direct geographical references, models must rely solely on provided context, enabling a more rigorous evaluation of their understanding of cultural norms and social reasoning.
-See Table 1 for examples.
 Country India Egypt Mexico
@@
 Table 1: Sample situations with contextualizations and labels from NormAd-Eti
+This approach enables us to generate situations across diverse cultural contexts and levels of norm adherence.
+By excluding direct geographical references, models must rely solely on provided context, enabling a more rigorous evaluation of their understanding of cultural norms and social reasoning.
+See Table 1 for examples.
 4.2 Automatic Filtration
@@
 These results indicate that the annotators overwhelmingly rated our situations, corresponding RoT s, Values, labels, and their relationships as valid, confirming the validity of NormAd.
-We report per-question scores and payments in Appendix B
-(a) RoT
-(b) Country +Value
-(c) Country
-Figure 3: Comparison of accuracies across LLaMa-1-SFT (7b, 13b, 30b), LLaMa-2 (7b, 13b, 70b), OLMo7b (SFT/Chat), GPT-3.5-turbo, GPT-4, and Mistral over the all three contexts.
-Models perform significantly worse in Country and Country +Value contexts compared to the RoT context.
-Baseline performance (no context) is reported in Appendix C and D.
+We report per-question scores and payments in Appendix B.
 4.4 Verification of Human Performance
@@
 This highlights that humans from the relevant culture show strong performance at determining the acceptability of situations when conditioned on abstract Value and Country contexts.
-See App.
-B.3 for country-wise splits.
+Please refer to Appendix B.3 for country-wise splits.
+(a) RoT
+(b) Country +Value
+(c) Country
+Figure 3: Comparison of accuracies across LLaMa-1-SFT (7b, 13b, 30b), LLaMa-2 (7b, 13b, 70b), OLMo7b (SFT/Chat), GPT-3.5-turbo, GPT-4, and Mistral over the all three contexts.
+Models perform significantly worse in Country and Country +Value contexts compared to the RoT context.
+Baseline performance (no context) is reported in Appendix C and D.
 5 Experimental Setup
+We evaluate several language models on their ability to adapt to varying levels of cultural contexts.
 5.1 Models
@@
 Value and Country
-LLMs show clear limitations when handling Country and Value +Country contexts, with the best performing models GPT-3.5-turbo, GPT-4footnote 6, and Mistral-7b-Instruct achieving only 59-63% accuracy for Value +Country and 51-56% for Country (see Figure 3).
+LLMs show clear limitations when handling Country and Value +Country contexts, with the best performing models GPT-3.5-turbo, GPT-4footnote 8, and Mistral-7b-Instruct achieving only 59-63% accuracy for Value +Country and 51-56% for Country (see Figure 3).
 In contrast, our human study across three countries (§4.4) demonstrates that humans can perform very well in these settings, achieving a high accuracy of 90%.
@@
 Our human study (§4.4) demonstrates that humans perform exceptionally well on this task, achieving high 95.6% accuracy.
-However, models under perform, likely due to a lack of adaptability to textual similarity under cultural and social nuances.
-The best performing models are GPT-4666 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance.
+However, models under perform, as shown in Figure 3, likely due to a lack of adaptability to cultural and social nuances in textual similarity tasks.
+The best performing models are GPT-4888 We note that our data was generated with GPT-4, which may give it an unfair advantage; however, even so, we find that GPT-4 still struggles with performance.
 with 87.6%, Mistral-7b-Instruct with 81.8% and Llama-2-70b-chat with 71.3%, lagging behind human performance.
@@
 (2024).
+Figure 4: Comparision of model accuracies under Country + Value shows a notable performance skew, with top models (with increased size or improved preference alignment methods) performing better in social situations from English-speaking countries than in African-Islamic cultural regions.
 6.2 How well do models perform across the Inglehart-Welzel (IW) cultural map?
 We mapped 75 countries into 8 clusters based on the Inglehart-Welzel cultural map.
-The Country + Value conditioned results, illustrated in Figure 4, show that best-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 6 vary in performance across different cultural zones.
+The Country + Value conditioned results, illustrated in Figure 4, show that best-performing models like Llama-2-70b, Llama-1-30b-SFT-KTO, and GPT-4footnote 8 vary in performance across different cultural zones.
 For instance, they perform better with situations from “English Speaking” countries (e.g., USA) than from “African-Islamic” countries (e.g., Saudi Arabia).
@@
 (2023).
-Figure 4: Model-wise accuracies under Country + Value show a notable performance skew, with top models performing better in social situations from English-speaking countries than in African-Islamic cultural zones.
 What is the effect of different preference alignment optimizations?
@@
 Figure 5 shows that KTO emerges as the most effective option for cultural adaptability, when conditioned on RoT.
-We see similar trends for Country and Value + Country as well (see App.
-D.1).
+We see similar trends for Country and Value + Country as well (see Appendix D.1 for more details).
+Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 8) across ground-truth labels.
+Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e.
+violations) and “neutral” (i.e irrelevant).
 6.3 What is the performance across subcategories of NormAd-Eti?
@@
 Further, gift-giving can be highly contextual in some cultures (Stauss, 2023), with differences in expense, presentation, and meaning playing a significant role in societal norms (Hanna and Srivastava, 2015).
+We additionally present our quantitative findings for subcategories in Figure 11 in Appendix D.
+Most models exhibit a performance dip for the ‘gifts’ subaxis.
+The Country + Value/ RoT contextualization mitigates this drop to some extent for some (but not all) models.
 This highlights the considerable adaptability required from LLMs in handling such complex social norms.
-Table 6 in App.
-D.4 presents some failure cases of GPT-3.5-turbo.
+Table 7 in Appendix D.4 presents some failure cases of GPT-3.5-turbo.
 6.4 How well do models do across social acceptabilities (Yes/No/Neutral)?
 We analyze how the social acceptability labels of situations affects model performance.
-Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6, Llama-1-30b-SFT-KTO).
+Figure 6 shows the averaged label-wise accuracies of our overall best-performing models (Llama-2-70b, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 8, Llama-1-30b-SFT-KTO).
 Models generally perform better on situations that conform to social norms (labeled ‘Yes’), and improve on norm-violating situations (labeled ‘No’) with increasing levels of context, indicating that inherent agreement biases within LLMs could impact their adaptability (Sun et al., 2024; Perez et al., 2022).
@@
 (2022).
-Figure 6: Averaged accuracy of best performing models (Llama-2-70b, Llama-1-30b-SFT-KTO, Mistral-Instruct, GPT-3.5-turbo, GPT-4footnote 6) across ground-truth labels.
-Models are biased towards “yes” (i.e conformations) and worse at “no” (i.e.
-violations) and “neutral” (i.e irrelevant).
 7 Conclusion
@@
 8 Limitations
-Our research examines LLM’s ability to adapt to cultural nuances through a test bed of social situations.
+Our research examines LLMs’ abilities to adapt to cultural nuances through a test bed of social situations.
 However, certain limitations present in our dataset and evaluation framework may warrant further research, such as:
 Existence of Multiple Cultural Proxies:
-Adilazuarda et al.
-(2024) acknowledges the difficulty of defining what ‘culture’ means in the context of language models, and categorize prior work according to cultural proxies, linguistic interactions and measurement strategies.
-NormAd can be thus categorized as a black-box LM evaluation employing a semantic proxy to test models’ reasoning abilities concerning etiquette-related social norms (§4.1), with analyses on demographically informed axes (§6.2).
-While our approach provides valuable insights on LLM’s cultural competencies, we believe that a more comprehensive picture of LLMs’ cultural landscape can be painted through other demographically and lingusitcally informed proxies and frameworks, and , by broadening the scope of evaluation Bhatt and Diaz (2024).
+Defining ‘culture’, especially in the context of language models is challenging, with prior work categorizing approaches by cultural proxies, linguistic interactions, and measurement strategies Adilazuarda et al.
+(2024).
+NormAd employs a black-box evaluation approach, using etiquette-related social norms as a semantic proxy of culture, with analyses on demographically informed axes (§6.2).
+While this approach offers valuable insights into LLMs’ cultural competencies, broader evaluation through diverse proxies is needed to capture the full cultural landscape (Bhatt and Diaz, 2024).
+Cultural Diversity and Representation:
+Cultural norms are highly diverse, with significant variation within countries, across regions, and among social classes.
+The Cultural Atlas only captures the dominant cultural narrative present in each country, leaving several variations unrepresented.
+Future work should should build resources that capture these diverse cultural perspectives and evaluate models on their ability to adapt across them.
 Multilingualism and Linguistic Variations:
-We conduct evaluations only in English, as Shen et al.
-(2024) highlight that prompting in English helps models leverage knowledge more effectively and mitigates issues arising from their varying linguistic capabilities and instruction-following skills.
-However, we acknowledge the significance of multilingualism and linguistic variations such as AAVE (African American Vernacular English) in ethical and cultural studies of AI (Li et al., 2024; Sap et al., 2019; Naous et al., 2023).
-Our study excludes these aspects from the evaluation due to the extensive validation efforts required for curating a synthetically generated and translated test bed across all languages and linguistic variations.
-This points to an area for future investigation for cultural adaptability studies, correlating multilingual reasoning performance with reasoning over cultures associated with such languages.
-Granularity of Cultural Norms:
-While our curated test bed encompasses narratives from 75 countries, it is essential to note the existence of more than 75 countries globally, each characterized by unique and finely-grained cultural norms.
-Additionally, there can be multiple cultural variations within a single country.
-We opted to scope our study based on the Cultural Atlas as a verified source of information, recognizing that certain nuances may not be fully captured within this framework.
+In this work, we conduct evaluations only in English.
+While testing across multiple languages and linguistic variations is essential for robust benchmarking of LLMs, it is beyond the scope of this study.
+Prior work highlights that prompting in English – given current skewed data representations – helps models leverage knowledge more effectively and mitigates issues arising from varying linguistic capabilities and instruction-following skills Shen et al.
+(2024).
+We encourage future work to investigate multilingual reasoning performance and its correlation with cultural adaptability across languages.
 Dynamic Cultural Evaluation
-Most current work in ‘culture’, including our own through NormAd-Eti, often treats the dynamic and multifaceted nature of culture as static variables during evaluations.
+As a pragmatic way for approaching culture, much research, including our own through NormAd-Eti, often treats the dynamic and multifaceted nature of culture as static variables during evaluation.
 This static approach may inadvertently perpetuate cultural stereotypes and fail to capture the continuously changing cultural nuances.
 To address these limitations, we suggest a modification to our evaluation framework, envisioned as future work, that would allow users to specify their own norms and values.
-Our framework, NormAd, is designed to be flexible, enabling these modifications easily, which is crucial for accommodating the evolving nature of cultural contexts.
-Improving our framework in this manner would greatly benefit the adaptability and cultural sensitivity of LLMs, ensuring they remain relevant and effective across diverse global environments.
+Our framework, NormAd, is designed to be flexible, which is crucial for accommodating the evolving nature of cultural contexts.
 Ethics Statement
-In this work, we study the cultural adaptability of LLMs – can LLMs align to human values across varying cultural contexts?
-While we suggest working on improving LLMs capabilities on this front, we acknowledge the existence of several human-computer interaction studies that assess the impact of personifying language models to cater to multiple demographics such as Black Americans (Harrington and Egede, 2023).
-These studies show that identity-related concepts such as age and racial likeness play a part in maintaining user trust and comfort.
-However, care must be taken when constructing personalized LLMs, owing to societal-risks such as the propagation of polarized views between historically conflicting demographics, and disparities in access among these subpopulations (Kirk et al., 2023).
-Finally, while we maintain equality over representing all demographics and cultures, we do not discuss the impact of the interaction and access disparities between these subpopulations on the use of language technologies, and the lack of representation they may cause.
+In this work, we study the cultural adaptability of LLMs – specifically, can LLMs align to human values across varying cultural contexts?
+While we advocate for improving LLM capabilities in this area, we recognize the complexities involved.
+Prior human-computer interaction studies suggest that personifying language models to cater to multiple demographics, such as Black Americans, can enhance user trust and comfort (Harrington and Egede, 2023; Wenzel and Kaufman, 2024).
+However, the extent to which LLMs should adapt to users’ cultural nuances remains uncertain.
+Excessive adaptation risks mimicry that may be perceived as manipulative, undermining user trust, particularly if the adaptation is seen as a shortcut to gaining social acceptance within a subgroup.
+Moreover, highly adaptable systems may amplify societal risks, such as reinforcing polarized views between historically conflicting demographics (Kirk et al., 2023).
+These complexities are further compounded by the fact that cultural norms are not monolithic; multiple variations often exist within a single country, region, or social group.
+Addressing this diversity requires adaptable frameworks that empower users to prescribe their cultural values, or opt out of certain adaptations altogether.
+Crucially, LLMs should adapt based on user-provided preferences rather than impose cultural norms.
+As a first step towards this, NormAd provides a framework for measuring cultural adaptability, with our benchmark NormAd-Eti merely serving as a proxy for measuring adaptability rather than prescribing cultural standards.
+Acknowledgements
+This project is funded in part by NSF grant (#2230466), DSO National Laboratories, and an Amazon Research Award, Spring 2023 CFP.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of Amazon.
+We additionally acknowledge the Cultural Atlas, a social cohesion project delivered by Mosaica and SBS, and the team that curated its cultural values, norms, and specifications.
+Finally, we would like to thank Shaily Bhatt and Joel Mire for their insightful comments on our work.
 Appendix A Appendix
@@
 Fleiss $\kappa$ and Krippendorff’s $\alpha$ are calculated.
+B.4 Column Mapping from the cultural atlas
+Original Column Mapped Column
+basic_etiquette basic_etiquette
+manners_in_vietnam basic_etiquette
+māori_etiquette basic_etiquette
+cleanliness basic_etiquette
+direct_manners basic_etiquette
+tipping basic_etiquette
+‘taarof’_(politeness_and_mutual_respect) basic_etiquette
+pub_etiquette basic_etiquette
+visiting visiting
+visiting_and_eating visiting
+visiting_a_village visiting
+eating eating
+eating_out eating
+religious_dietary_laws eating
+drinking eating
+drinking_coffee eating
+toasting eating
+gifts gifts
+gift-giving gifts
+gift_giving gifts
+offering_and_complimenting_items gifts
+Table 5: Mapping of Original Columns to Mapped Columns
 Appendix C F1-scores over NormAd-Eti across all models
@@
 (blue represents country, yellow represents value, green represents rule-of-thumb.
+Dashed line represents baseline performance with no conditioning.
+(a) Archangel_7b_sft
+(b) Archangel_7b_sft_kto
+(c) Archangel_7b_sft_ppo
+(d) Archangel_7b_sft_dpo
+(e) Archangel_13b_sft
+(f) Archangel_13b_sft_ppo
+(g) Archangel_13b_sft_dpo
+(h) Archangel_13b_sft_kto
+(i) Archangel_30b_sft
+(j) Archangel_30b_sft_ppo
+(k) Archangel_30b_sft_dpo
+(l) Archangel_30b_sft_kto
+(m) llama2-7b-chat
+(n) llama2-13b-chat
+(o) llama2-70b-chat
+(p) olmo-7b-sft
+(q) olmo-7b-instruct
+(r) gpt-3.5-turbo-0125
+(s) gpt4
+(t) mistral-chat
+Figure 11: Accuracy across subaxes (eating, visiting, gifts, basic_etiquette) for all contextualizations across all models.
+Blue represents country, yellow represents country+value, green represents rule-of-thumb.
 Dashed line represents baseline performance with no conditioning.